### v1

Abstract.
For a positive integer $k$, the $k$-cut complex of a graph $G$ is the simplicial complex whose facets are the $(|V(G)|-k)$-subsets $\sigma$ of the vertex set $V(G)$ of $G$ such that the induced subgraph of $G$ on $V(G)\setminus\sigma$ is disconnected. These complexes first appeared in the master thesis of Denker and were further studied by Bayer et al. in [Topology of cut complexes of graphs, SIAM Journal on Discrete Mathematics, 2024]. In the same article, Bayer et al. conjectured that for $k\geq 3$, the $k$-cut complexes of squared cycle graphs are shellable. Moreover, they also conjectured about the Betti numbers of these complexes when $k=3$. In this article, we prove these conjectures for $k=3$.
2020 Mathematics Subject Classification:
57M15, 52B22, 55U05, 05C69, 05E45
1. Introduction
All the graphs in this article are assumed to be finite and simple, that is, without loops and multiple edges. We denote the set of vertices and the set of edges of a graph $G$ by $V(G)$ and $E(G)$, respectively. For an integer $k\geq 1$, the $k$-cut complex of a graph $G$, denoted as $\Delta_{k}(G)$, is the simplicial complex whose facets (maximal simplices) are $\sigma\subseteq V(G)$ such that $|\sigma|=|V(G)|-k$ and the induced subgraph $G[V(G)\setminus\sigma]$ is disconnected. These complexes first appeared in [5] and were further studied by Bayer et al. in [1]. One of the main motivations behind cut complexes was a famous theorem of Ralf Fröberg [7] connecting commutative algebra and graph theory through topology (see Theorem 1.1).
Let $\Delta$ be a simplicial complex on the vertex set $V(\Delta)=\{v_{1},v_{2},\ldots,v_{n}\}$ and let $\mathbb{K}$ be a field. The Stanley-Reisner ideal $I_{\Delta}$ of $\Delta$ is the ideal of the polynomial ring $\mathbb{K}[x_{1},\ldots,x_{n}]$ generated by the monomials corresponding to minimal subsets of $V(\Delta)$, which are not simplices of $\Delta$, i.e., $I_{\Delta}=\langle x_{i_{1}}\ldots x_{i_{k}}:\{v_{i_{1}},v_{i_{2}},\ldots,v_{i_{k}}\}\notin\Delta\rangle$. The Stanley-Reisner ring $\mathbb{K}[\Delta]$ is the quotient ring $\mathbb{K}[x_{1},\ldots,x_{n}]/I_{\Delta}$. For more details, we refer the reader to [6].
The Alexander dual $\Delta^{\vee}$ of the simplicial complex $\Delta$ is the simplicial complex on the vertex set $V(\Delta)$, whose simplices are the subsets of $V(\Delta)$ such that their complements are not simplices of $\Delta$, i.e.,
$\Delta^{\vee}=\{\sigma\subset V(\Delta):V(\Delta)\setminus\sigma\notin\Delta\}.$
The clique complex $\mathsf{Cl}(G)$ of a graph $G$ is the simplicial complex whose simplices are $\sigma\subseteq V(G)$ such that the induced subgraph $G[\sigma]$ is a complete graph. It is easy to check that the Stanley-Reisner ideal $I_{\mathsf{Cl}(G)}$ of $\mathsf{Cl}(G)$, is generated by quadratic square-free monomials.
Theorem 1.1 ([7, Theorem 1], [6, p. 274]).
A Stanley–Reisner ideal $I_{\Delta}$ generated by quadratic square-free monomials has a $2$-linear resolution if and only if $\Delta$ is the clique complex $\mathsf{Cl}(G)$ of a chordal graph $G$.
Theorem 1.2 ([6, Proposition 8]).
The following are equivalent for a graph $G$.
  1. (1)
    $G$ is chordal.
  2. (2)
    $\mathsf{Cl}(G)^{\vee}$ is Cohen-Macaulay over a field $k$.
  3. (3)
    $\mathsf{Cl}(G)^{\vee}$ is vertex decomposable.
For a pure simplicial complex, it is well known (see [3, Section 11]) that
[…]
Observe that $\mathsf{Cl}(G)^{\vee}=\Delta_{2}(G)$ for any graph $G$. Therefore, Theorem 1.2 implies the following.
(1.1) $G\ \text{is chordal}\Longleftrightarrow\Delta_{2}(G)\ \text{is shellable}\Longleftrightarrow\Delta_{2}(G)\ \text{is vertex decomposable}.$
For definitions of vertex decomposable, Cohen-Macaulay, and shellable complexes, see Section 2. For more details on these concepts, we refer to [3, Section 11] and [9, Section 3.6].
Inspired from Fröberg’s theorem (Theorem 1.1) and (1.1), Denker [5] introduced $k$-cut complexes of graphs which are a generalization of $\Delta_{2}(G)=\mathsf{Cl}(G)^{\vee}$. In [2], Bayer et al. introduced an another generalization of $\Delta_{2}(G)$, the total $k$-cut complexes. For $k\geq 1$, the total $k$-cut complex of a graph $G$, denoted as $\Delta_{k}^{t}(G)$, is the simplicial complex whose facets are $\sigma\subseteq V(G)$ such that $|\sigma|=|V(G)|-k$ and the induced subgraph $G[V(G)\setminus\sigma]$ does not contain any edge. In particular, $\Delta_{2}(G)=\Delta_{2}^{t}(G)$ for any graph $G$.
For $n\geq 3$, the squared cycle graph $W_{n}$ is a graph where the set of vertices $V(W_{n})=\{0,1,2,\ldots,n-1\}$ and the set of edges $E(W_{n})=\{\{i,i+1\ (\text{mod}\ n)\},\{i,i+2\ (\text{mod}\ n)\}:0\leq i\leq n-1\}$. For $n\geq 5$, squared cycle graphs $W_{n}$ are also $4$-regular circulant graphs and hence Cayley graphs of the cyclic group on $n$ elements (for the definition of circulant graphs see Section 2). In [2, Proposition 4.19], authors studied the total $2$-cut complexes of squared cycle graphs and proved that these complexes are homotopy equivalent to wedge of spheres of dimension lower than the dimension of $\Delta_{2}^{t}(W_{n})=\Delta_{2}(W_{n})$. Therefore, the complexes $\Delta_{2}(W_{n})$ are not shellable. Further, based on their observations supported by Sage calculations (for $n\leq 13$ and $k\leq 5$), they conjectured that $\Delta_{k}(W_{n})$ are shellable for $k\geq 3$.
Conjecture 1.3 ([1, Conjecture 7.25]).
For $k\geq 3$, the cut complex $\Delta_{k}(W_{n})$ is shellable for $n\geq k+6$. For $k=3$ and $n\geq 9,$ the Betti numbers are $\binom{n-4}{2}-9=\{1,6,12,19,27,\dots\}$.
In this article, we prove 1.3 for $k=3$. More precisely, the main result of this article is the following.
Theorem 1.4.
Let $n\geq 9$. Then the $3$-cut complex $\Delta_{3}(W_{n})$ of a squared cycle graph is shellable. Moreover,
$\Delta_{3}(W_{n})\simeq\bigvee\limits_{\binom{n-4}{2}-9}\mathbb{S}^{n-4}.$
Note that the sequence $\binom{n-4}{2}-9=\{1,6,12,19,27,\dots\}$ is OEIS A051936.
This paper is organized as follows: In Section 2, we provide the necessary preliminaries related to graph theory and simplicial complexes. In Section 3, we give the proof of Theorem 1.4. This section is divided into two subsections. To prove Theorem 1.4, we define an order in the facets of $\Delta_{3}(W_{n})$ and in Section 3.1, we show that this order is a shelling order. In Section 3.2, we characterize and count the number of spanning facets for the shelling order to conclude the number of spheres appearing in the wedge in the homotopy type of $\Delta_{3}(W_{n})$.
2. Preliminaries
In this section, we recall some basic definitions and results used in this article.
2.1. Graph
A graph $G$ is a pair $(V(G),E(G))$, where $V(G)$ is the set of vertices of $G$ and $E(G)\subset\binom{V(G)}{2}$ denotes the set of edges. If $\{x,y\}\in E(G)$, it is also denoted by $x\sim y$ and we say that $x$ is adjacent to $y$. A subgraph $H$ of $G$ is a graph with $V(H)\subset V(G)$ and $E(H)\subset E(G)$. For a subset $U\subseteq V(G)$, the induced subgraph $G[U]$ is the subgraph whose set of vertices is $V(G[U])=U$ and the set of edges is $E(G[U])=\{\{a,b\}\in E(G)\ |\ a,b\in U\}$.
A graph is connected if there exists a path between each pair of vertices. A graph is disconnected if it is not connected. The graph with the empty set $\emptyset$ as its set of vertices is considered connected.
Let $n\geq 2$ be a positive integer and $S\subset\{1,2,\ldots,n-1\}$. The circulant graph $C_{n}(S)$ is the graph whose set of vertices is $V(C_{n}(S))=\{0,1,2,\ldots,n-1\}$ and any two vertices $x$ and $y$ are adjacent if and only if $x-y\ (\text{mod \ $n$})\in S\cup-S$, where $-S=\{n-a\ |\ a\in S\}$. Circulant graphs are also Cayley graphs of $\mathbb{Z}_{n}$, the cyclic group on $n$ elements. For $n\geq 3,$ the squared cycle graph $W_{n}$ on $n$ vertices, is a circulant graph $C_{n}(S),$ where $S=\{1,2\}.$
We refer the reader to [4] and [11] for more details about the graphs.
2.2. Simplicial complex
A finite abstract simplicial complex $\Delta$ is a collection of finite sets such that if $\tau\in\Delta$ and $\sigma\subset\tau$, then $\sigma\in\Delta$. The elements of $\Delta$ are called simplices of $\Delta$. The dimension of a simplex $\sigma$ is equal to $|\sigma|-1$. The dimension of an abstract simplicial complex is the maximum of the dimensions of its simplices. The 00-dimensional simplices are called vertices of $\Delta$, and the set of vertices of $\Delta$ is denoted by $V(\Delta)$. If $\sigma\subset\tau$, we say that $\sigma$ is a face of $\tau$. If a simplex has dimension $d$, it is said to be $d$-dimensional. A simplex that is not a face of any other simplex is called a maximal simplex or facet. The set of maximal simplices of $\Delta$ is denoted by $M(\Delta)$. A simplicial complex is called pure $d$-dimensional, if all of its maximal simplices are of dimension $d$.
In this article, we consider any simplicial complex as a topological space, namely, its geometric realization. For the definition of geometric realization, we refer to the book [10] by Kozlov. For terminologies of algebraic topology used in this article, see [8].
Let $\Delta$ be a simplicial complex and let $\tau$ be a simplex of $\Delta$. The link of $\tau$ is the simplicial complex defined as
$lk_{\Delta}(\tau):=\{\sigma\in\Delta\ |\ \sigma\cap\tau=\emptyset,\text{ and }\sigma\cup\tau\in\Delta\}.$
The deletion of $\tau$ is the simplicial complex defined as
$dl_{\Delta}(\tau):=\{\sigma\in\Delta\ |\ \tau\not\subseteq\sigma\}.$
Definition 2.1.
A pure simplicial complex $\Delta$ is called vertex decomposable if it is empty or if there exists a vertex $v$ of $\Delta$ such that both $lk_{\Delta}(v)$ and $dl_{\Delta}(v)$ are vertex decomposable.
Definition 2.2.
A complex $\Delta$ is $p$-acyclic over a field $\mathbb{K}$ if the reduced homology group $\tilde{H}_{i}(\Delta;\mathbb{K})$ vanishes for $i\leq p$.
Definition 2.3.
Let $\Delta$ be a pure simplicial complex and $\mathbb{K}$ be a field. Then $\Delta$ is Cohen-Macaulay over $\mathbb{K}$ if $lk_{\Delta}(\sigma)$ is $(dim\ lk_{\Delta}(\sigma)-1)$-acyclic for each $\sigma$ in $\Delta$.
2.3. Shellability
Definition 2.4.
[…]
In other words, a simplicial complex $\Delta$ has a shelling order $F_{1},F_{2},\ldots,F_{t}$ of its facets if and only if for any $i,j$ satisfying $1\leq i<j\leq t$, there exists $1\leq r<j$ such that $F_{r}\cap F_{j}=F_{j}\setminus\{\lambda\}$ for some $\lambda\in F_{j}\setminus F_{i}$.
Definition 2.5 ([10, Section 12.1]).
Let $\Delta$ be a shellable simplicial complex, and let $F_{1},F_{2},\ldots,F_{t}$ be a shelling on $\Delta$. Then $F_{j}$ $(1<j\leq t)$ is said to be a spanning facet for the given shelling order if for each $\lambda\in F_{j}$, there exists $r<j$ such that $F_{r}\cap F_{j}=F_{j}\setminus\{\lambda\}$.
Theorem 2.6.
[10, Theorem 12.3] Let $\Delta$ be a pure shellable simplicial complex of dimension $d$. Then $\Delta$ has the homotopy type of a wedge of $\beta$ spheres of dimension $d$, where $\beta$ is the number of total spanning facets in a given shelling. Hence
$\Delta\simeq\bigvee\limits_{\beta}\mathbb{S}^{d}.$
3. Proof of Theorem 1.4
In this section, we prove the main result (Theorem 1.4) of this article. Throughout the section, we fix $n\geq 9$. Our aim is to define an order $\prec$ in the facets of $\Delta_{3}(W_{n})$, which provides a shelling order.
Let
[…]
We arrange the elements of $V(W_{n})=\{0,1,2,\ldots,n-1\}$ into an ordered set $\mathcal{O}:=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$, where $\alpha_{t}=m+(-1)^{t-1}\lfloor(t/2)\rfloor\ (\text{mod}\ n)$ $\forall$ $1\leq t\leq n$ and $\lfloor(t/2)\rfloor$ denotes the greatest integer less than or equal to $t/2$. Hence
[…]
We define an order $<_{\mathcal{O}}$ in the elements of $V(W_{n})$ as follows: For any $x,y\in V(W_{n})$, we say that $x<_{\mathcal{O}}y$ if and only if $x=\alpha_{s}$ and $y=\alpha_{t}$ for some $1\leq s<t\leq n$.
Remark 3.1.
Let $x,y\in V(W_{n})$.
  1. (i)
    If $x<m$, then $x<_{\mathcal{O}}y$ if and only if either $y<x$ or $y\geq 2m-x$.
  2. (ii)
    If $x\geq m$, then $x<_{\mathcal{O}}y$ if and only if either $y<2m-x$ or $y>x$.
  3. (iii)
    If $y<m$, then $x<_{\mathcal{O}}y$ if and only if $y<x<2m-y$.
  4. (iv)
    If $y>m$, then $x<_{\mathcal{O}}y$ if and only if $2m-y\leq x<y$.
Throughout this section, we denote the complement of a set $X\subseteq V(W_{n})$ by $X^{c}$.
Let $\mathcal{T}$ be the set of all $(n-3)$-subsets of $V(W_{n})$. We partition $\mathcal{T}$ into disjoint sets $\mathcal{T}_{s}$ as follows: Let $\mathcal{T}_{1}:=\{\tau\in\mathcal{T}:\alpha_{1}\in\tau^{c}\}.$ For $s>1$, define
$\mathcal{T}_{s}:=\{\tau\in\mathcal{T}:\alpha_{s}\in\tau^{c}\ \text{and }\alpha_{1},\alpha_{2},\dots,\alpha_{s-1}\notin\tau^{c}\}.$
Clearly $\mathcal{T}=\bigsqcup\mathcal{T}_{s}$.
Throughout the section, if $\tau\in\mathcal{T}$ such that $\tau\in\mathcal{T}_{s}$ for some $s\geq 1,$ and $\tau^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$, then we assume that $s_{1}<s_{2}$.
Remark 3.2.
Let $\tau\in\mathcal{T}$ be such that $\tau\in\mathcal{T}_{s}$ for some $s\geq 1,$ and ${\tau}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Observe that $\alpha_{s}<_{\mathcal{O}}s_{1},s_{2}$ by definition of $\mathcal{T}_{s}$.
Let $M(\Delta_{3}(W_{n}))$ be the set of facets of $\Delta_{3}(W_{n})$. For any $F\in M(\Delta_{3}(W_{n}))$, we have $|F|=n-3$ and $W_{n}[F^{c}]$ is disconnected. Clearly, $M(\Delta_{3}(W_{n}))\subset\mathcal{T}=\bigsqcup\mathcal{T}_{s}$.
Remark 3.3.
Let $F\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Observe that since $W_{n}[F^{c}]$ is disconnected and $\alpha_{s}<_{\mathcal{O}}s_{1},s_{2}$, we get $\alpha_{s}\notin\{0,1,n-1\}$. Hence $22\leq\alpha_{s}\leq n-$ 2.
For each $s$, we define an order $<_{\mathcal{T}_{s}}$ in the elements of $\mathcal{T}_{s}$ as follows: if $F$, $F^{\prime}$in $\mathcal{T}_{s}$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s^{\prime}_{1},s^{\prime}_{2}\}$, then $F<_{\mathcal{T}_{s}}F^{\prime}$ if and only if either $s_{1}<s^{\prime}_{1}$ or, $s_{1}=s^{\prime}_{1}$ and $s_{2}<s^{\prime}_{2}$.
Now using the order $<_{\mathcal{T}_{s}}$, we define an order $\ll$ in the elements of $M(\Delta_{3}(W_{n}))$ as follows: Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$. Then $F\ll F^{\prime}$ if and only if either $s<t$ or, $s=t$ and $F<_{\mathcal{T}_{s}}F^{\prime}$.
To get our desired shelling order $\prec$, we slightly modify the order $\ll$ by “displacing” some elements of the poset $(M(\Delta_{3}(W_{n})),\ll)$ and get a new poset $(M(\Delta_{3}(W_{n})),$ $\prec$)≺ ). For this, we first define a subset $\mathcal{D}$ of $M(\Delta_{3}(W_{n}))$ as follows: A facet $F\in\mathcal{D}$ if and only if there exists $s$ such that $F\in\mathcal{T}_{s}$ and $F$ satisfies any of the following conditions:
  1. $(\mathcal{D}_{1})$
    ${F}^{c}=\{\alpha_{s}\}\sqcup\{\alpha_{s}-3,\alpha_{s}+1\}$, where $\alpha_{s}=m+1$.
  2. $(\mathcal{D}_{2})$
    ${F}^{c}=\{\alpha_{s}\}\sqcup\{\alpha_{s}-1,\alpha_{s}+3\}$, where $\alpha_{s}=m-1$.
  3. $(\mathcal{D}_{3})$
    ${F}^{c}=\{\alpha_{s}\}\sqcup\{\alpha_{s}-4,\alpha_{s}-3\}$, where $\alpha_{s}=m+1$.
  4. $(\mathcal{D}_{4})$
    ${F}^{c}=\{\alpha_{s}\}\sqcup\{\alpha_{s}+3,\alpha_{s}+4\}$, where $\alpha_{s}\in\{m-1,m,\dots,n-6\}$.
Definition 3.4.
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. The order $\prec$ on $M(\Delta_{3}(W_{n}))$ is defined as follows: We say $F\prec F^{\prime}$ if and only if any one of the following conditions is true:
  1. (i)
    $F,F^{\prime}\notin\mathcal{D}$ and $F\ll F^{\prime}$.
  2. (ii)
    $F\in\mathcal{D}$, $F^{\prime}\notin\mathcal{D}$ such that $F\in\mathcal{T}_{s}$, $F^{\prime}\in\mathcal{T}_{t}$ for some $s<t$.
  3. (iii)
    $F\notin\mathcal{D}$, $F^{\prime}\in\mathcal{D}$ such that $F\in\mathcal{T}_{s}$, $F^{\prime}\in\mathcal{T}_{t}$, and either $s=t$ or $s<t$ for some $s$ and $t$.
  4. (iv)
    $F,F^{\prime}\in\mathcal{D}$ and $F\ll F^{\prime}$.
If none of these conditions (i)-(iv) is true, we say $F^{\prime}\prec F$.
Proposition 3.5.
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$ for some $s\neq t$ such that ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and ${F^{\prime}}^{c}=({F}^{c}\setminus\{\nu\})\sqcup\{\lambda\}$ for some $\nu\in{F}^{c}$ and $\lambda\in F$. If $\lambda<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$.
Proof.
Suppose $\nu=\alpha_{s}$. By Remark 3.2, $\nu<_{\mathcal{O}}s_{1},s_{2}$. Since $\lambda<_{\mathcal{O}}\nu$, $\lambda<_{\mathcal{O}}s_{1},s_{2}$. Thus $F^{\prime}\in\mathcal{T}_{t}$ implies that $F^{\prime c}=\{\alpha_{t}\}\sqcup\{s_{1},s_{2}\}$, where $\alpha_{t}=\lambda$ and we get $t<s$ by the definition of $<_{\mathcal{O}}$. Therefore, $F^{\prime}\prec F$ by Definition 3.4.
Now suppose $\nu\in\{s_{1},s_{2}\}$. First, let $\nu=s_{1}$. Then $F^{\prime c}=\{\lambda,\alpha_{s},s_{2}\}$. Since $\lambda<_{\mathcal{O}}\alpha_{s}$ and $F^{\prime c}\in\mathcal{T}_{t}$, either ${F^{\prime}}^{c}=\{\alpha_{t}\}\sqcup\{\alpha_{s},s_{2}\}$ or ${F^{\prime}}^{c}=\{\alpha_{t}\}\sqcup\{s_{2},\alpha_{s}\}$, where $\alpha_{t}=\lambda$. Therefore $t<s$ by the definition of $<_{\mathcal{O}}$, and hence, $F^{\prime}\prec F$ using Definition 3.4. If $\nu=s_{2}$, we get $F^{\prime}\prec F$ using a similar argument. ∎
To prove Theorem 1.4, in Section 3.1, we first show that the order $\prec$ given in Definition 3.4 provides a shelling order for $\Delta_{3}(W_{n})$. Later, in Section 3.2, we identify and count all the spanning facets for this shelling order.
3.1. Shelling Order
In this section, we prove that the order $\prec$ given in Definition 3.4 provides a shelling order for $\Delta_{3}(W_{n})$. To demonstrate that $\prec$ provides a shelling order for $\Delta_{3}(W_{n})$, by definition of shellability, we need to prove the following: For any $F_{i},F_{j}\in M(\Delta_{3}(W_{n}))$ with $F_{i}\prec F_{j}$, there is a facet $F_{r}$ such that
($\ast$) $F_{r}\prec F_{j}\text{ and }F_{r}\cap F_{j}=F_{j}\setminus\{\lambda\}\text{ for some }\lambda\in F_{j}\setminus F_{i}.\hfill$
Proposition 3.6.
Let $F_{i}\prec F_{j}$ be two elements of $M(\Delta_{3}(W_{n}))$. If there is an $F\in M(\Delta_{3}(W_{n}))$ such that $F\prec F_{j}$ and $F^{c}=(F_{j}^{c}\setminus\{\nu\})\sqcup\{\lambda\}$ for some $\nu\in F_{j}^{c}$ and $\lambda\in F_{j}\setminus F_{i}$, then $F_{r}=F$ satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
Proof.
We already have $F\prec F_{j}$. Since $F^{c}=(F_{j}^{c}\setminus\{\nu\})\sqcup\{\lambda\}$, we get $F=(F_{j}\setminus\{\lambda\})\sqcup\{\nu\}$ and thus, $F\cap F_{j}=F_{j}\setminus\{\lambda\}$, where $\lambda\in F_{j}\setminus F_{i}$. Hence $F_{r}=F$ satisfies ($\ast$3.1) for the pair $F_{i}\prec F_{j}$. ∎
Let $F_{i},F_{j}\in M(\Delta_{3}(W_{n}))$ such that $F_{i}\prec F_{j}$. Let $F_{i}\in\mathcal{T}_{i_{0}}$ and $F_{j}\in\mathcal{T}_{j_{0}}$, where ${F_{i}}^{c}=\{\alpha_{i_{0}}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j_{0}}\}\sqcup\{j_{1},j_{2}\}$. By our assumption, we have $i_{1}<i_{2}$ and $j_{1}<j_{2}$. For convenience, we drop the subscripts $i_{0}$ and $j_{0}$ from $\alpha_{i_{0}}$ and $\alpha_{j_{0}}$, and refer to them simply as $\alpha_{i}$ and $\alpha_{j}$, respectively. Our aim is to find an $F_{r}$ which satisfies ($\ast$3.1) for the pair $F_{i}\prec F_{j}$.
We deal with the cases $F_{j}\in\mathcal{D}$ and $F_{j}\notin\mathcal{D}$ separately. The case $F_{j}\in\mathcal{D}$ is considered in Lemma 3.9, and the case $F_{j}\notin\mathcal{D}$ is considered in Lemmas 3.7 and 3.8. In Lemma 3.7, we deal with the case $i_{0}=j_{0}$, and in Lemma 3.8, we deal with the case $i_{0}\neq j_{0}$.
Lemma 3.7.
Suppose that $F_{j}\notin\mathcal{D}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then there exists an $F_{r}\in M(\Delta_{3}(W_{n}))$ that satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
Proof.
Since $F_{i}\in\mathcal{T}_{j_{0}}$, we have $\alpha_{i}=\alpha_{j}$, and thus ${F_{i}}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$. If $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}\neq\emptyset$, then ($\ast$3.1) is satisfied by taking $F_{r}=F_{i}$. So, we assume that $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$.
Since $F_{i}\prec F_{j}$, $F_{j}\notin\mathcal{D}$ and $F_{i},F_{j}\in\mathcal{T}_{j_{0}}$, using Definition 3.4, we get $F_{i}\notin\mathcal{D}$ and $F_{i}\ll F_{j}$. Further, since $F_{i},F_{j}\in\mathcal{T}_{j_{0}}$, $F_{i}\ll F_{j}$ implies that $F_{i}<_{\mathcal{T}_{j_{0}}}F_{j}$. This means that either $i_{1}<j_{1}$, or $i_{1}=j_{1}$ and $i_{2}<j_{2}$. Since $i_{1},i_{2},j_{1}$, and $j_{2}$ are distinct, we get $i_{1}<j_{1}<j_{2}$. Moreover, $i_{1}<i_{2}$. By Remark 3.3, we have $22\leq\alpha_{j}\leq n-$ 2.
We now consider the following cases: (I) $j_{1},j_{2}<\alpha_{j}$, (II) $j_{1},j_{2}>\alpha_{j}$ and (III) $j_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}$.
  1. (I)
    $j_{1},j_{2}<\alpha_{j}$.
    Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Since $i_{1}<j_{1}<j_{2}<\alpha_{j}$, we have $i_{1}<\alpha_{j}-2$. Suppose $i_{1}\sim\alpha_{j}$. Then $\alpha_{j}\leq n-2$ implies that $i_{1}=0$ and $\alpha_{j}=n-2$. Since $W_{n}[{F_{i}}^{c}]$ is disconnected, we get $i_{2}\neq n-1$ Therefore, $i_{1},i_{2},j_{1},j_{2}<\alpha_{j}$. We have $\alpha_{j}=n-2>m$ as $n\geq 9$) and $\alpha_{j}<_{\mathcal{O}}i_{1},i_{2},j_{1},j_{2}$. Thus, by Remark 3.1 (ii), $i_{1},i_{2},j_{1},j_{2}<2m-\alpha_{j}=2m-n+2\leq 3$, a contradiction as $i_{1},i_{2},j_{1},j_{2}$ are distinct. Hence $i_{1}\nsim\alpha_{j}.$ Since $W_{n}[{F_{j}}^{c}]$ is disconnected, we have $j_{2}\nsim\alpha_{j}$ or $j_{2}\nsim j_{1}$. If $j_{2}\nsim\alpha_{j}$, then $\alpha_{j}$ is an isolated vertex in $W_{n}[F^{\prime c}]$; and if $j_{2}\nsim j_{1}$, then $j_{2}\nsim i_{1}$, which implies that $i_{1}$ is an isolated vertex in $W_{n}[F^{\prime c}]$. This means that $W_{n}[{F^{\prime}}^{c}]$ is disconnected and hence $F^{\prime}\in M\Delta_{3}(W_{n}))$.
    Since $i_{1}<j_{1}<j_{2}<\alpha_{j}$, we observe that $F^{\prime}$ does not satisfy any of the conditions $(\mathcal{D}_{1})$-$(\mathcal{D}_{4})$, and therefore $F^{\prime}\not\in\mathcal{D}$. We have $F^{\prime}\in\mathcal{T}_{j_{0}}$. So, $i_{1}<j_{1}$ implies that $F^{\prime}<_{\mathcal{T}_{j_{0}}}F$ and thus $F^{\prime}\ll F_{j}$. Hence $F^{\prime}\prec F_{j}$ by Definition 3.4 (i), and thus, $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
  2. (II)
    $j_{1},j_{2}>\alpha_{j}$.
    Since $W_{n}[{F_{j}}^{c}]$ is disconnected, we get $j_{2}>\alpha_{j}+3$. Thus, $\alpha_{j}\geq 2$ implies that $j_{2}\nsim\alpha_{j}$. Further, $F_{i}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},i_{2}\}$. So we have either $i_{1}<\alpha_{j}$, or $i_{1}>\alpha_{j}$.
    1. (a)
      $i_{1}<\alpha_{j}$.
      Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$. We show that if $F^{\prime}\in M\Delta_{3}(W_{n}))$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1), otherwise $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). First, assume that $F^{\prime}\in M\Delta_{3}(W_{n}))$. Using the facts that $i_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}+3$, we see that $F^{\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$. Therefore $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
      We now assume that $F^{\prime}\notin M\Delta_{3}(W_{n}))$ i.e., $W_{n}[F^{\prime c}]$ is connected. Since $j_{2}\nsim\alpha_{j}$, we get $\alpha_{j}\sim i_{1}$ and $i_{1}\sim j_{2}.$ Thus, $j_{2}>j_{1}>\alpha_{j}>i_{1}$ implies that $i_{1}\geq\alpha_{j}-2$. We know that $i_{1}<i_{2}$ and $W_{n}[F_{i}^{c}]$ is disconnected. So $i_{2}>\alpha_{j}$. Since $i_{1},i_{2},j_{1}$ and $j_{2}$ are distinct, it follows that $i_{2}\leq n-3$ or $j_{1}\leq n-3$.
      We have $i_{2},j_{1}>\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}i_{2},j_{1}$. If $\alpha_{j}<m$, then using Remark 3.1 (i), $i_{2},j_{1}\geq 2m-\alpha_{j}$; and if $\alpha_{j}>m$, then $i_{2},j_{1}>\alpha_{j}>2m-\alpha_{j}$. Hence we conclude that $i_{2},j_{1}\geq 2m-\alpha_{j}$. This means that $n-3\geq 2m-\alpha_{j}\geq n-\alpha_{j}$. Hence $\alpha_{j}\geq 3$. Now $i_{1}\geq\alpha_{j}-2$ and $i_{1}\sim j_{2}$ imply that $\alpha_{j}=3$, $i_{1}=1$ and $j_{2}=n-1.$ If $2m=n+1,$ then $2m-\alpha_{j}=n-2>n-3,$ a contradiction. So $2m=n.$ Then $32m-\alpha_{j}=n-$ 3, and this implies that $j_{1}\in\{n-3,n-2\}$. Hence $j_{1}\geq 6$ (as $n\geq 9$). Therefore, since $i_{1}=1$ and $\alpha_{j}=3$, it follows that $j_{1}\nsim i_{1}$ and $j_{1}\nsim\alpha_{j}$. Clearly, $W_{n}[F^{\prime\prime c}]$ is disconnected i.e., $F^{\prime\prime}\in M\Delta_{3}(W_{n}))$. Now, using the facts that $i_{1}=\alpha_{j}-2$ and $j_{1}\geq\alpha_{j}+3$, we have $F^{\prime\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime\prime}\ll F_{j}$. Thus $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (i). Hence $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    2. (b)
      $i_{1}>\alpha_{j}$.
      Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Suppose $W_{n}[{F^{\prime}}^{c}]$ is connected. Then $2\leq\alpha_{j}<i_{1}<j_{1}<j_{2}$ implies that $i_{1}\leq\alpha_{j}+2$, $j_{1}=i_{1}+1$ and $j_{2}=i_{1}+2$. If $i_{1}=\alpha_{j}+1$, then $j_{1}=\alpha_{j}+2$ and $j_{2}=\alpha_{j}+3$, which implies that $W_{n}[{F_{j}}^{c}]$ is connected, a contradiction. So $i_{1}=\alpha_{j}+2$. This means that $j_{1}=\alpha_{j}+3$ and $j_{2}=\alpha_{j}+4$. Since $i_{1}<i_{2}$ and $i_{1},i_{2},j_{1},j_{2}$ are distinct, $\alpha_{j}\leq n-6$. Further, by Remark 3.1 (i) and (ii), $i_{1}>\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}i_{1}$ implies that $i_{1}\geq 2m-\alpha_{j}$ i.e., $\alpha_{j}+2\geq 2m-\alpha_{j}$. Hence, $\alpha_{j}\geq m-1$. This implies that $F_{j}$ satisfies $(\mathcal{D}_{4})$, which contradicts our assumption that $F_{j}\notin\mathcal{D}$. Therefore $W_{n}[{F^{\prime}}^{c}]$ is disconnected and hence $F^{\prime}\in M\Delta_{3}(W_{n}))$.
      Since $\alpha_{j}<i_{1}<j_{1}<j_{2}$, we have $F^{\prime}\notin\mathcal{D}$. Further, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, and thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
  3. (III)
    $j_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}$.
    Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$, ${F^{\prime\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$ and $F^{\prime\prime\prime c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$.
    First, let $F^{\prime\prime}\in M\Delta_{3}(W_{n}))$. Since $i_{1}<j_{1}$, we have $F^{\prime\prime}\ll F$. Therefore, if $F^{\prime\prime}\notin\mathcal{D}$, then $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (i). Hence, by taking $F_{r}=F^{\prime\prime}$, ($\ast$3.1) is satisfied.
    Suppose $F^{\prime\prime}\in\mathcal{D}$. Since $i_{1}<j_{1}<\alpha_{j}$, it follows that $F^{\prime\prime}$ satisfies $(\mathcal{D}_{3})$. This means that $i_{1}=\alpha_{j}-4$, $j_{1}=\alpha_{j}-3$, and $\alpha_{j}=m+1$. Note that $n\geq 9$ implies $m\geq 5$. So $\alpha_{j}\geq 6$. Now, $i_{1}=\alpha_{j}-4\geq 2$ and $j_{2}>\alpha_{j}$ implies that $i_{1}\nsim\alpha_{j}$ and $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$ and $F^{\prime}\notin\mathcal{D}$. Also, since $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Therefore, taking $F_{r}=F^{\prime}$, ($\ast$3.1) is satisfied by Proposition 3.6.
    Now, let $F^{\prime\prime}\notin M(\Delta_{3}(W_{n}))$, i.e., $W_{n}[F^{\prime\prime c}]$ is connected. We have $i_{1}<j_{1}<\alpha_{j}$ and $22\leq\alpha_{j}\leq n-$ 2. Suppose $j_{1}\nsim i_{1}$ or $j_{1}\nsim\alpha_{j}$. Then $i_{1}\sim\alpha_{j}$ such that $i_{1}=0$ and $\alpha_{j}=n-2$. Therefore, $\alpha_{j}<_{\mathcal{O}}i_{2}$ implies that $i_{2}\in\{1,2,n-1\}$ by definition of $\mathcal{O}$. We get a contradiction to the fact that $W_{n}[F_{i}^{c}]$ is disconnected.
    Thus, $j_{1}\sim i_{1}$ and $j_{1}\sim\alpha_{j}$. Now, since $W_{n}[F_{j}^{c}]$ is disconnected, $j_{2}\nsim\alpha_{j}$, which implies that $j_{2}\geq\alpha_{j}+3$. We have ${F^{\prime}}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Let $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Since $i_{1}<j_{1}<\alpha_{j}$ and $j_{2}\geq\alpha_{j}+3$, $F^{\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, and thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Hence, by taking $F_{r}=F^{\prime}$, ($\ast$3.1) is satisfied.
    We now assume that $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, i.e., $W_{n}[F^{\prime c}]$ is connected. Since $j_{2}\nsim\alpha_{j}$, we get $j_{2}\sim i_{1}$ and $i_{1}\sim\alpha_{j}$. Then $i_{1}<j_{1}<\alpha_{j}<j_{2}$ implies that $i_{1}=\alpha_{j}-2$ and $j_{1}=\alpha_{j}-1$. Also, $j_{2}\sim i_{1}$ implies that either $j_{2}=n-2$ and $i_{1}=0$, or $j_{2}=n-1$ and $i_{1}\in\{0,1\}$. If $i_{1}=0$, then $\alpha_{j}=2$, and thus, $\alpha_{j}<_{\mathcal{O}}i_{2}$ implies that $i_{2}\in\{1,n-2,n-1\}$ by definition of $\mathcal{O}$. This means that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. So, $i_{1}=1$, $j_{1}=2$, $\alpha_{j}=3$ and $j_{2}=n-1$. Further, $W_{n}[F_{i}^{c}]$ is disconnected implies that $i_{2}\nsim\alpha_{j}$, and since $i_{2}\neq j_{2}$, $5<i_{2}<n-1=j_{2}$. It follows that $i_{2}\nsim j_{1}$. We have ${F^{\prime\prime\prime}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Clearly, $W_{n}[F^{\prime\prime\prime c}]\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-1$, $i_{2}\geq\alpha_{j}+3$ and $\alpha_{j}=3\leq m-2$ (as $m\geq 5$), we have $F^{\prime\prime\prime}\notin\mathcal{D}$. Also, $i_{2}<j_{2}$ implies that $F^{\prime\prime\prime}\ll F_{j}$, and thus $F^{\prime\prime\prime}\prec F_{j}$ by Definition 3.4 (i). Hence, $F_{r}=F^{\prime\prime\prime}$ satisfies ($\ast$3.1).
Lemma 3.8.
[…]
Proof.
We have ${F_{i}}^{c}=\{\alpha_{i}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$ such that $F_{i}\prec F_{j}$. Then $F_{j}\notin\mathcal{D}$ and $F_{i}\notin\mathcal{T}_{j_{0}}$ implies that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ by Definition 3.4.
Let ${F_{q}}^{c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. Observe that if $F_{q}\in M(\Delta_{3}(W_{n}))$, then $F_{q}\prec F_{j}$ by Proposition 3.5, and hence $F_{r}=F_{q}$ satisfies ($\ast$3.1) by Proposition 3.6.
Clearly, $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ implies that $\alpha_{j}\neq m$. Thus, we have two cases: (A) $\alpha_{j}<m$, and (B) $\alpha_{j}>m$.
Case A: $\alpha_{j}<m$.
Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, we get $\alpha_{j}<\alpha_{i}<2m-\alpha_{j}$ by Remark 3.1 (iii). We have $2m-\alpha_{j}>m$. It follows that either $\alpha_{j}<\alpha_{i}<m$, or $m\leq\alpha_{i}<2m-\alpha_{j}$.
Subcase A.1: $\alpha_{j}<\alpha_{i}<m$.
In this case, we show that $F_{q}\in M(\Delta_{3}(W_{n}))$, and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1).
  1. (I)
    $j_{1},j_{2}<\alpha_{j}$.
    We have $j_{1}<j_{2}<\alpha_{j}<\alpha_{i}<m<n-1$. Thus $j_{1}\nsim\alpha_{i}$. Also, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. It follows that $F_{q}\in M(\Delta_{3}(W_{n}))$.
  2. (II)
    $j_{1},j_{2}>\alpha_{j}$.
    Since $\alpha_{j}<\alpha_{i}<m$, $\alpha_{j}\leq m-2$. Further, $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}>\alpha_{j}$ implies that $j_{1}\geq m+2$ by definition of $\mathcal{O}$. We have $\alpha_{j}<\alpha_{i}<m<m+2\leq j_{1}<j_{2}$. Thus $j_{1}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$.
  3. (III)
    $j_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}$.
    Since $\alpha_{j}<\alpha_{i}<m$, $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}>\alpha_{j}$, we get $j_{2}\geq m+2$. We have $j_{1}<\alpha_{j}<\alpha_{i}<m<m+2\leq j_{2}$. Thus $j_{2}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
Subcase A.2: $m\leq\alpha_{i}<2m-\alpha_{j}$.
In this case, we show that either $F_{q}\in M(\Delta_{3}(W_{n}))$ and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1), or $F_{q}\notin M(\Delta_{3}(W_{n}))$ and there exists some $F_{r}\neq F_{q}$ that satisfies ($\ast$3.1) for the pair $F_{i}\prec F_{j}$.
  1. (I)
    $j_{1},j_{2}<\alpha_{j}$.
    Observe that since $j_{1},j_{2}<\alpha_{j}$, if $\alpha_{j}<4$, then $W_{n}[F_{j}^{c}]$ is connected. So $\alpha_{j}\geq 4$, and hence $\alpha_{i}<2m-\alpha_{j}\leq 2m-4\leq n-3$. We have $j_{1}<j_{2}<\alpha_{j}<m\leq\alpha_{i}<n-3$. Thus $j_{1}\nsim\alpha_{i}$. Since $W_{n}[F_{j}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$.
  2. (II)
    $j_{1},j_{2}>\alpha_{j}$.
    Using the facts that $j_{1},j_{2}>\alpha_{j}$, $\alpha_{j}<m$, and $\alpha_{j}<_{\mathcal{O}}j_{1},j_{2}$, it follows that $j_{1},j_{2}\geq 2m-\alpha_{j}>\alpha_{i}$ by Remark 3.1 (i). Thus $\alpha_{j}<m\leq\alpha_{i}<j_{1}<j_{2}$.
    Observe that if $j_{1}\nsim\alpha_{i}$ or $j_{1}\nsim j_{2}$, then $\alpha_{i}\geq m$ implies that $j_{2}\nsim\alpha_{i}$, and hence $F_{q}\in M(\Delta_{3}(W_{n}))$. So assume that $j_{1}\sim\alpha_{i}$ and $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{\alpha_{j},j_{2}\}$. We first show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$.
    Since $W_{n}[F_{j}^{c}]$ is disconnected, $j_{1}\sim j_{2}$ implies that $j_{1}\nsim\alpha_{j}$ and $j_{2}\nsim\alpha_{j}$. Therefore, if $\alpha_{i}\nsim\alpha_{j}$ or $\alpha_{i}\nsim j_{2}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $\alpha_{i}\sim\alpha_{j}$. We need to show that $\alpha_{i}\nsim j_{2}$. Since $\alpha_{j}<m$ and $\alpha_{i}\geq m$, $\alpha_{i}\sim\alpha_{j}$ implies that $\alpha_{j}\in\{m-2,m-1\}$. If $\alpha_{j}=m-2$, then $\alpha_{i}=m$ and $j_{2}>j_{1}\geq m+2$ (as $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}>\alpha_{j}$). Thus $\alpha_{i}\nsim j_{2}$.
    If $\alpha_{j}=m-1$, then $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ implies that $\alpha_{i}=m$. Further, we have $j_{1}\sim\alpha_{i}$, $j_{1}\nsim\alpha_{j}$, and $j_{1}>\alpha_{i}=m$. Thus $j_{1}=m+2$, and hence $j_{2}>m+2$. This means that $\alpha_{i}\nsim j_{2}$.
    Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, from Proposition 3.5, we see that $F^{\prime}\prec F_{j}$. Thus $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
  3. (III)
    $j_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}$.
    We have $j_{2}>\alpha_{j}$, $\alpha_{j}<m$, and $\alpha_{j}<_{\mathcal{O}}j_{2}$. Thus, we get $j_{2}\geq 2m-\alpha_{j}>\alpha_{i}$ by Remark 3.1 (i). So, $j_{1}<\alpha_{j}<m\leq\alpha_{i}<j_{2}$.
    Suppose $j_{2}\nsim\alpha_{i}$. Now, since $W_{n}[{F_{j}}^{c}]$ is disconnected, we have $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{2}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{1}\nsim\alpha_{j}$, then $j_{1}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. Thus, we see that $F_{q}\in M(\Delta_{3}(W_{n}))$.
    Now, we assume that $j_{2}\sim\alpha_{i}$. Suppose $j_{1}\sim\alpha_{i}$. Then, $j_{1}<\alpha_{j}<m\leq\alpha_{i}<j_{2}$ implies that either $j_{1}=m-2$, $\alpha_{j}=m-1$, $\alpha_{i}=m$ and $j_{2}\in\{m+1,m+2\}$, or $j_{1}=0$, $\alpha_{i}=n-2$, and $j_{2}=n-1$. In the former case, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{2}=m+2=\alpha_{j}+3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{2})$, a contradiction as $F_{j}\notin\mathcal{D}$. Hence, we have $j_{1}=0$, $\alpha_{i}=n-2$, and $j_{2}=n-1$. Now, $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}<m$ implies that $\alpha_{j}\leq 2$. It follows that $W_{n}[{F_{j}}^{c}]$ is connected, again a contradiction. Therefore, $j_{1}\nsim\alpha_{i}$. Now, if $j_{1}\nsim j_{2}$, then $F_{q}\in M(\Delta_{3}(W_{n}))$.
    So, let $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{j_{1},\alpha_{j}\}$. Since $W_{n}[F_{j}^{c}]$ is disconnected and $j_{1}\sim j_{2}$, we get $j_{1}\nsim\alpha_{j}$. Therefore, $j_{1}\nsim\alpha_{i}$ implies that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. We have $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$. Thus $F^{\prime}\prec F_{j}$ by Proposition 3.5. Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
Case B: $\alpha_{j}>m$.
We have ${F_{q}}^{c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. Recall that, if $F_{q}\in M(\Delta_{3}(W_{n}))$, then $F_{r}=F_{q}$ satisfies ($\ast$3.1). Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, $2m-\alpha_{j}\leq\alpha_{i}<\alpha_{j}$ by Remark 3.1 (iv). Further, $2m-\alpha_{j}<m$. This means that either $2m-\alpha_{j}\leq\alpha_{i}\leq m$, or $m<\alpha_{i}<\alpha_{j}$.
Subcase B.1: $2m-\alpha_{j}\leq\alpha_{i}\leq m$.
In this case also, we show that either $F_{q}\in M(\Delta_{3}(W_{n}))$, or $F_{q}\notin M(\Delta_{3}(W_{n}))$ and we find some $F_{r}\neq F_{q}$ that satisfies ($\ast$3.1).
  1. (I)
    $j_{1},j_{2}<\alpha_{j}$.
    Since $\alpha_{j}<_{\mathcal{O}}j_{1},j_{2}$ and $\alpha_{j}>m$, it follows from Remark 3.1 (ii) that $j_{1},j_{2}<2m-\alpha_{j}\leq\alpha_{i}$. So, we have $j_{1}<j_{2}<\alpha_{i}\leq m<\alpha_{j}$.
    Observe that if $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{i}$, then $\alpha_{i}\leq m$ implies that $j_{1}\nsim\alpha_{i}$, and hence $F_{q}\in M(\Delta_{3}(W_{n}))$. So assume that $j_{1}\sim j_{2}$ and $j_{2}\sim\alpha_{i}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let ${F^{\prime}}^{c}:=\{\alpha_{i}\}\sqcup\{j_{1},\alpha_{j}\}$. We first show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$.
    Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\sim j_{2}$ implies that $j_{1}\not\sim\alpha_{j}$. Therefore, if $\alpha_{i}\nsim\alpha_{j}$ or $\alpha_{i}\nsim j_{1}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $\alpha_{i}\sim\alpha_{j}$. We show that $\alpha_{i}\nsim j_{1}$. Suppose $\alpha_{i}\sim j_{1}$. Since $\alpha_{i}\leq m$ and $\alpha_{j}>m$, $\alpha_{i}\sim\alpha_{j}$ implies that $\alpha_{j}\in\{m+1,m+2\}$. If $\alpha_{j}=m+1$, then $\alpha_{i}\in\{m-1,m\}$. Using the facts that $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}<\alpha_{j}$, it follows that $j_{2}\leq m-2$. Hence $j_{1}<m-2$. Then $\alpha_{i}=m-1$, $j_{1}=m-3=\alpha_{j}-4$ and $j_{2}=m-2=\alpha_{j}-3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{3})$, a contradiction.
    If $\alpha_{j}=m+2$, then $\alpha_{i}=m$ and $j_{1}<j_{2}\leq m-3$ (as $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}<\alpha_{j}$). Thus $\alpha_{i}\nsim j_{1}$, which contradicts our assumption that $\alpha_{i}\sim j_{1}$. Since we get a contradiction in each case, our assumption that $\alpha_{i}\sim j_{1}$ is false. Therefore, $\alpha_{i}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$.
    Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, $F\prec F_{j}$ by Proposition 3.5. Thus $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
  2. (II)
    $j_{1},j_{2}>\alpha_{j}$.
    If $\alpha_{j}\geq n-4$, then $j_{1},j_{2}>\alpha_{j}$ implies $W_{n}[F_{j}^{c}]$ is connected. So $\alpha_{j}\leq n-5$, and hence $\alpha_{i}\geq 2m-\alpha_{j}\geq 5$. We have $j_{2}>j_{1}>\alpha_{j}>m\geq\alpha_{i}\geq 5$. Thus $j_{2}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
  3. (III)
    $j_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}$.
    Since $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $\alpha_{j}>m$, we get $j_{1}<2m-\alpha_{j}\leq\alpha_{i}$ by Remark 3.1 (ii). So, we have $j_{1}<\alpha_{i}\leq m<\alpha_{j}<j_{2}$.
    Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. Now, if $j_{1}\nsim\alpha_{i}$, then $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{i}$ and $j_{2}\nsim\alpha_{j}$ imply that $j_{2}\nsim\alpha_{i}$).
    We now assume that $j_{1}\sim\alpha_{i}$. Suppose $j_{2}\sim\alpha_{i}$. Then $j_{1}<\alpha_{i}\leq m<\alpha_{j}<j_{2}$ implies that either $\alpha_{i}=m$, $\alpha_{j}=m+1$, $j_{2}=m+2$ and $j_{1}\in\{m-2,m-1\}$, or $\alpha_{i}=1$, $j_{2}=n-1$, and $j_{1}=0$. In the former case, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}=m-2=\alpha_{j}-3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{1})$, a contradiction. So, we have $\alpha_{i}=1$, $j_{2}=n-1$, and $j_{1}=0$. Then $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}>m$ implies that $\alpha_{j}=n-1=j_{2}$, which is again a contradiction. Therefore, $j_{2}\nsim\alpha_{i}$. Now, if $j_{1}\nsim j_{2}$, then $F_{q}\in M(\Delta_{3}(W_{n}))$.
    So, let $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{\alpha_{j},j_{2}\}$. Since $W_{n}[F_{j}^{c}]$ is disconnected and $j_{1}\sim j_{2}$, we get $j_{2}\nsim\alpha_{j}$. Therefore, $j_{2}\nsim\alpha_{i}$ implies that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, we get $F^{\prime}\prec F_{j}$ by Proposition 3.5. Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
Subcase B.2: $m<\alpha_{i}<\alpha_{j}$.
We show that $F_{q}\in M(\Delta_{3}(W_{n}))$, and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1).
  1. (I)
    $j_{1},j_{2}<\alpha_{j}$.
    Since $m<\alpha_{i}<\alpha_{j}$, we have $\alpha_{j}\geq m+2$. Then $\alpha_{j}<_{\mathcal{O}}j_{2}<\alpha_{j}$ implies that $j_{2}\leq m-3$. It follows that $j_{1}<j_{2}\leq m-3<m<\alpha_{i}<\alpha_{j}$. Thus $j_{2}\nsim\alpha_{i}$. Further, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. This implies that $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
  2. (II)
    $j_{1},j_{2}>\alpha_{j}$.
    We have $j_{2}>j_{1}>\alpha_{j}>\alpha_{i}>m>1$. So $j_{2}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
  3. (III)
    $j_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}$.
    Since $m<\alpha_{i}<\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}j_{1}<\alpha_{j}$, we get $j_{1}\leq m-3$. Hence $j_{1}\leq m-3<m<\alpha_{i}<\alpha_{j}<j_{2}$. Thus $j_{1}\nsim\alpha_{i}$. Now, since $W_{n}[{F_{j}}^{c}]$ is disconnected, we have $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. This implies that $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{2}\nsim\alpha_{j}$ implies that $j_{2}\nsim\alpha_{i}$).
In the following Lemma, we now consider the case when $F_{j}\in\mathcal{D}$.
Lemma 3.9.
Suppose that $F_{j}\in\mathcal{D}$. Then there exists an $F_{r}\in M(\Delta_{3}(W_{n}))$ that satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
Proof.
We have ${F_{i}}^{c}=\{\alpha_{i}\}\sqcup\{i_{1},i_{2}\}$, ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$ and $F_{i}\prec F_{j}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, i.e., $i_{0}=j_{0}$, then we have $\alpha_{i}=\alpha_{j}$. In this case, if $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}\neq\emptyset$, then $F_{r}=F_{i}$ satisfies ($\ast$3.1). So we assume that $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$ whenever $F_{i}\in\mathcal{T}_{j_{0}}$. Moreover, if $F_{i}\notin\mathcal{T}_{j_{0}}$, then $F_{i}\prec F_{j}$ and $F_{j}\in\mathcal{D}$ imply that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ by Definition 3.4. Note that since $n\geq 9$, we have $45\leq m\leq n-$ 4. Now, $F_{j}\in\mathcal{D}$ implies that $F_{j}$ satisfies one of the conditions from $(\mathcal{D}_{1})$ to $(\mathcal{D}_{4})$.
  1. (1)
    $F_{j}$ satisfies $(\mathcal{D}_{1})$.
    We have ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{\alpha_{j}-3,\alpha_{j}+1\}$ and $\alpha_{j}=m+1$. Hence, $36\leq\alpha_{j}\leq n-$ 3.
    Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. First we show that if $i_{1}<j_{1}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>j_{2}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq j_{1}$ and $i_{2}\leq j_{2}$.
    First, suppose $i_{1}<j_{1}$. Since $i_{1}<j_{1}=\alpha_{j}-3<\alpha_{j}\leq n-3$, we get $\alpha_{j}\nsim i_{1}$ and $\alpha_{j}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$. Therefore, $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii) and (iv). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Let $i_{2}>j_{2}$. Since $\alpha_{j}\geq 6$, $j_{1}=\alpha_{j}-3\geq 3$. So we have $3\leq j_{1}=\alpha_{j}-3<\alpha_{j}<\alpha_{j}+1=j_{2}<i_{2}$. It follows that $j_{1}\nsim\alpha_{j}$ and $j_{1}\nsim i_{2}$. Hence $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-3$ and $i_{2}\geq\alpha_{j}+2$, $F^{\prime\prime}\notin\mathcal{D}$. Therefore $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (iii), and thus $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Now, suppose $i_{1}\geq j_{1}$ and $i_{2}\leq j_{2}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then $\alpha_{j}=\alpha_{i}\neq i_{2}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{2}\leq j_{2}=\alpha_{j}+1$ implies that $i_{2}<\alpha_{j}$. Therefore, we have $\alpha_{j}-3=j_{1}<i_{1}<i_{2}<\alpha_{j}$. It follows that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
    We know that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m+1$. Thus, $\alpha_{i}\in\{m-1,m\}$. We have $m-2=j_{1}\leq i_{1}<i_{2}\leq j_{2}=m+2$. Therefore, since $W_{n}[F_{i}^{c}]$ is disconnected, we get $\alpha_{i}=m-1$, $i_{1}=j_{1}=m-2$ and $i_{2}=j_{2}=m+2$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
  2. (2)
    $F_{j}$ satisfies $(\mathcal{D}_{2})$.
    We have ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{\alpha_{j}-1,\alpha_{j}+3\}$, where $\alpha_{j}=m-1$. Therefore $45\leq m\leq n-$ 4 implies that $54\leq\alpha_{j}\leq n-$ 5.
    Suppose $j_{1}\leq i_{1}<i_{2}\leq j_{2}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then we have $\alpha_{j}=\alpha_{i}\neq i_{1}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. This implies that $\alpha_{j}<i_{1}<i_{2}<j_{2}=\alpha_{j}+3$. It follows that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. So $F_{i}\notin\mathcal{T}_{j_{0}}$.
    We have $m-2=j_{1}\leq i_{1}<i_{2}\leq j_{2}=m+2$. Moreover, since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m-1$, we get $\alpha_{i}=m$. It follows that $W_{n}[F_{i}^{c}]$ is connected, again a contradiction. Hence our assumption that $j_{1}\leq i_{1}<i_{2}\leq j_{2}$ is false.
    Now, we consider two cases: $i_{1}<j_{1}$ and $i_{1}\geq j_{1}$.
    Suppose $i_{1}<j_{1}$. Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. First, we show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. We have $i_{1}<j_{1}=\alpha_{j}-1<\alpha_{j}<\alpha_{j}+3=j_{2}$. So $j_{2}\nsim\alpha_{j}$. Now, if $i_{1}\nsim\alpha_{j}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. So, let $i_{1}\sim\alpha_{j}$. Then $i_{1}=\alpha_{j}-2\geq 2$ (as $\alpha_{j}\geq 4$). Thus $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $i_{1}<j_{1}=\alpha_{j}-1$, $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Now, suppose that $i_{1}\geq j_{1}$. If $i_{2}\leq j_{2}$, then $j_{1}\leq i_{1}<i_{2}\leq j_{2}$, which is a contradiction. So $i_{2}>j_{2}$. Let $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Since $\alpha_{j}\geq 4$, we have $j_{1}\geq 3$. Then $3\leq j_{1}=\alpha_{j}-1<\alpha_{j}<\alpha_{j}+3=j_{2}<i_{2}$. It follows that $i_{2}\nsim j_{1}$ and $i_{2}\nsim\alpha_{j}$. Hence, $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-1$ and $i_{2}\geq\alpha_{j}+4$, $F^{\prime\prime}\notin\mathcal{D}$. Thus $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (iii) and $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1).
  3. (3)
    $F_{j}$ satisfies $(\mathcal{D}_{3})$.
    We have ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{\alpha_{j}-4,\alpha_{j}-3\}$, where $\alpha_{j}=m+1$. Therefore $36\leq\alpha_{j}\leq n-$ 3 (as $45\leq m\leq n-$ 4).
    Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. First we show that if $i_{1}<j_{1}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>\alpha_{j}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq j_{1}$ and $i_{2}\leq\alpha_{j}$.
    Suppose $i_{1}<j_{1}$. Since $i_{1}<j_{1}=\alpha_{j}-4<\alpha_{j}\leq n-3$, we get $\alpha_{j}\nsim i_{1}$ and $\alpha_{j}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $i_{1}<j_{1}=\alpha_{j}-4$, $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Let $i_{2}>\alpha_{j}$. Since $\alpha_{j}\geq 6$, we have $2\leq j_{1}=\alpha_{j}-4<\alpha_{j}<i_{2}$. So, $j_{1}\nsim\alpha_{j}$ and $j_{1}\nsim i_{2}$. Thus $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-4$ and $i_{2}\geq\alpha_{j}+1$, it follows that $F^{\prime\prime}\notin\mathcal{D}$. Hence $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1).
    Now, let $i_{1}\geq j_{1}$ and $i_{2}\leq\alpha_{j}$. Suppose $F_{i}\in\mathcal{T}_{j_{0}}$. Then $\alpha_{j}=\alpha_{i}\neq i_{2}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{1}\geq j_{1}=j_{2}-1$ implies that $i_{1}>j_{2}$. So, we have $\alpha_{j}-3=j_{2}<i_{1}<i_{2}<\alpha_{j}$. It follows that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
    Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m+1$, we get $\alpha_{i}\in\{m-1,m\}$. We have $m-3=j_{1}\leq i_{1}<i_{2}\leq\alpha_{j}=m+1$. Therefore, $W_{n}[F_{i}^{c}]$ is disconnected implies that $\alpha_{i}=m$, $i_{1}=j_{1}=m-3$ and $i_{2}=\alpha_{j}=m+1$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
  4. (4)
    $F_{j}$ satisfies $(\mathcal{D}_{4})$.
    We have ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{\alpha_{j}+3,\alpha_{j}+4\}$, where $\alpha_{j}\in\{m-1,m,\dots,n-6\}$. Since $m\geq 5$, we have $64\leq\alpha_{j}\leq n-$ 6.
    Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{2},i_{2}\}$. First we show that if $i_{1}<\alpha_{j}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>j_{2}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq\alpha_{j}$ and $i_{2}\leq j_{2}$.
    Suppose $i_{1}<\alpha_{j}$. We first show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{j}\leq n-6$, we get $j_{2}=\alpha_{j}+4\leq n-2$. We have $i_{1}<\alpha_{j}<\alpha_{j}+4=j_{2}\leq n-2$. So $j_{2}\nsim\alpha_{j}$. Now, if $i_{1}\nsim\alpha_{j}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. So, let $i_{1}\sim\alpha_{j}$. Then $i_{1}\in\{\alpha_{j}-2,\alpha_{j}-1\}$ and therefore $i_{1}\geq 2$. Thus $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $i_{1}\leq\alpha_{j}-1$ and $j_{2}=\alpha_{j}+4$, it follows that $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
    Now, suppose $i_{2}>j_{2}$. We have $4\leq\alpha_{j}<\alpha_{j}+4=j_{2}<i_{2}$. Thus $\alpha_{j}\nsim j_{2}$ and $\alpha_{j}\nsim i_{2}$. Hence, $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $i_{2}>j_{2}=\alpha_{j}+4$, $F^{\prime\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
    Finally, let $i_{1}\geq\alpha_{j}$ and $i_{2}\leq j_{2}$. Suppose $F_{i}\in\mathcal{T}_{j_{0}}$. Then $\alpha_{j}=\alpha_{i}\neq i_{1}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{2}\leq j_{2}=j_{1}+1$ implies that $i_{2}<j_{1}$. Therefore, we have $\alpha_{j}<i_{1}<i_{2}<j_{1}=\alpha_{j}+3$. This implies that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
    We have $\alpha_{j}\in\{m-1,m,\ldots,n-6\}$. Since $F_{i}\notin\mathcal{T}_{j_{0}}$ implies that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, we have $\alpha_{j}\neq m$. We consider two cases: (i) $\alpha_{j}=m-1$, and (ii) $\alpha_{j}\in\{m+1,m+2,m+3,\ldots,n-6\}$.
    1. (i)
      $\alpha_{j}=m-1$.
      Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, $\alpha_{i}=m$. We have $m-1=\alpha_{j}\leq i_{1}<i_{2}\leq j_{2}=\alpha_{j}+4=m+3$. Therefore, $W_{n}[F_{i}^{c}]$ is disconnected implies that $i_{1}=\alpha_{j}=m-1$ and $i_{2}=j_{2}=m+3$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
    2. (ii)
      $\alpha_{j}\in\{m+1,m+2,m+3,\ldots,n-6\}$.
      Let $F^{\prime\prime\prime c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. We have $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}>m$. So by Remark 3.1 (iv), $\alpha_{j}>\alpha_{i}\geq 2m-\alpha_{j}\geq 2m-(n-6)\geq 6$. Then, we have $6\leq\alpha_{i}<\alpha_{j}<\alpha_{j}+3=j_{1}<j_{2}$. It follows that $\alpha_{i}\nsim j_{1}$ and $\alpha_{i}\nsim j_{2}$. Hence $F^{\prime\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, Proposition 3.5 implies that $F^{\prime\prime\prime}\prec F_{j}$. Hence $F_{r}=F^{\prime\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
From Lemmas 3.7, 3.8 and 3.9, it follows that for any $F_{i},F_{j}\in M(\Delta_{3}(W_{n}))$ with $F_{i}\prec F_{j}$, there exists an $F_{r}\prec F_{j}$ such that $F_{i}\cap F_{j}\subseteq F_{r}\cap F_{j}$ and $|F_{r}\cap F_{j}|=$ $|F_{j}|-1$. This means that $\prec$ provides a shelling order for $\Delta_{3}(W_{n})$.
3.2. Spanning Facets
In this section, we characterize and count the spanning facets for the shelling order $\prec$.
Remark 3.10.
Observe that by definition, any facet $F\in M(\Delta_{3}(W_{n}))$ is a spanning facet if and only if for each $\mu\in F$, there exists $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$.
Define a subset $\mathcal{S}$ of $M(\Delta_{3}(W_{n}))$ as follows: A facet $F\in\mathcal{S}$ if and only if there exists $s$ such that $F\in\mathcal{T}_{s}$ and $F$ satisfies one of the following conditions:
  1. $(\mathcal{S}_{1})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}=3$ and $s_{1}\in V(W_{n})\setminus(\{0,1,2,3,\ldots,2m-4\}\sqcup\{n-1\})$.
  2. $(\mathcal{S}_{2})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{4,5,\ldots,m-2\}$ and $s_{1}\in V(W_{n})\setminus(\{\alpha_{s}-4,\alpha_{s}-3,\alpha_{s}-2\}\sqcup\{\alpha_{s},\alpha_{s}+1,\ldots,2m-\alpha_{s}-1\}\sqcup\{n-1\})$.
  3. $(\mathcal{S}_{3})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{m-1,m,m+1,m+2,\ldots,n-3\}$, $s_{1}\in V(W_{n})\setminus(\{\omega,\omega+1,\ldots,\alpha_{s},\alpha_{s}+1,\alpha_{s}+2,\alpha_{s}+3\text{ (mod $n$})\}\sqcup\{y\})$, $\omega=\min\{2m-\alpha_{s},\alpha_{s}-4\}$ and
    $y=\begin{cases}n-1&\text{ if }\alpha_{s}<n-4,\\ 0&\text{ if }\alpha_{s}=n-4,\\ 1&\text{ if }\alpha_{s}>n-4.\end{cases}$
Our aim is to show that a facet $F\in M(\Delta_{3}(W_{n}))$ is a spanning facet if and only if $F\in\mathcal{S}$.
Proposition 3.11.
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $F\in\mathcal{T}_{s}$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$ and ${F^{\prime}}^{c}=\{\alpha_{s},s_{1},\mu\}$ for some $\mu\in F$. If $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\ll F$.
Proof.
Since $\alpha_{s}<_{\mathcal{O}}\mu,s_{1}$, we have $F^{\prime}\in\mathcal{T}_{s}$. If $\mu<s_{1}$, then $F^{\prime c}=\{\alpha_{s}\}\sqcup\{\mu,s_{1}\}$ and $F^{\prime}\ll F$. On the other hand, if $\mu>s_{1}$, then $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s_{1},\mu\}$, and thus $\mu<n-1$ implies that $F^{\prime}\ll F$. ∎
Proposition 3.12.
Let $F\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$. Let $\mathcal{U}=V(W_{n})\setminus\bigl{(}\{\alpha_{s}-4\ (\text{mod $n$}),\alpha_{s}-3,\alpha_{s}-2,\ldots,2m-\alpha_{s}-1\}\cup\{n-1\}\bigr{)}$. Suppose for each $\alpha_{s}\in\{3,4,\ldots,m-2\}$, $s_{1}\in\mathcal{U}$. Then $F$ is a spanning facet.
Proof.
From Remark 3.10, it is sufficient to show that for each $\mu\in F$, there exists a facet $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$.
Let $\mu\in F$ and $F^{\prime c}:=(F^{c}\setminus\{n-1\})\sqcup\{\mu\}=\{\alpha_{s},s_{1},\mu\}$. Since $\alpha_{s}\leq m-2$, we have $32m-\alpha_{s}-1\geq m+1\geq\alpha_{s}+$ 3, and hence $32m-\alpha_{s}-1\geq\alpha_{s}+$ 3. First, suppose that $s_{1}\neq\alpha_{s}+4$. Then $s_{1}\in\mathcal{U}$ implies that $s_{1}\notin\{\alpha_{s}-4\text{ (mod $n$)},\alpha_{s}-3,\alpha_{s}-2,\ldots,\alpha_{s}+4\}\cup\{n-1\}$. This means that $\alpha_{s}\nsim s_{1}$ and $|\alpha_{s}-s_{1}|\geq 5$. Further, since $\alpha_{s}\geq 3$ and $s_{1}<n-1$, we get $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Thus $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. So, let $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\ll F$ by Proposition 3.11. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Hence, $F^{\prime}\prec F$ by Definition 3.4 (i). Thus, we take $\tilde{F}=F^{\prime}$.
Now suppose that $s_{1}=\alpha_{s}+4$. First, let $\mu\neq\alpha_{s}+2$. We have $13\leq\alpha_{s}<\alpha_{s}+4=s_{1}<n-$ 1. It follows that $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Further, since $\alpha_{s}\nsim s_{1}$, we get $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. So assume that $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\ll F$ by Proposition 3.11. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Therefore, $F^{\prime}\prec F$ by Definition 3.4 (i), and we take $\tilde{F}=F^{\prime}$.
Now, let $\mu=\alpha_{s}+2$. Let $F^{\prime\prime c}:=(F^{c}\setminus\{s_{1}\})\sqcup\{\mu\}=\{\alpha_{s},\mu,n-1\}$. Since $13\leq\alpha_{s}<\alpha_{s}+2=\mu=s_{1}-2<s_{1}<n-$ 1, it follows that $\alpha_{s}\nsim n-1$ and $\mu\nsim n-1$. Therefore, $W_{n}[F^{\prime\prime c}]$ is disconnected. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime\prime}\prec F$ by Proposition 3.5 and we take $\tilde{F}=F^{\prime\prime}$. So, let $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime\prime}\in\mathcal{T}_{s}$ and thus, $F^{\prime\prime c}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. Since $\mu=\alpha_{s}+2<\alpha_{s}+4=s_{1}$, we have $F^{\prime\prime}\ll F$. Also, $\alpha_{s}<m-1$ implies that $F,F^{\prime\prime}\notin\mathcal{D}$. Hence, $F^{\prime\prime}\prec F$ by Definition 3.4 (i) and we take $\tilde{F}=F^{\prime\prime}$.
Therefore, in each case, we have a facet $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Thus, $F$ is a spanning facet. ∎
Lemma 3.13.
Suppose that $F\in\mathcal{S}$. Then $F$ is a spanning facet.
Proof.
Since $F\in\mathcal{S}$, we have $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$. Let $\mathcal{U}=V(W_{n})\setminus\bigl{(}\{\alpha_{s}-4\ (\text{mod $n$}),\alpha_{s}-3,\alpha_{s}-2,\ldots,2m-\alpha_{s}-1\}\cup\{n-1\}\bigr{)}$.
Let $\mu\in F$. Using Remark 3.10, it is sufficient to find a facet $\tilde{F}$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Let $F^{\prime c}:=(F^{c}\setminus\{n-1\})\sqcup\{\mu\}=\{\alpha_{s},s_{1},\mu\}$ and $F^{\prime\prime c}:=(F^{c}\setminus\{s_{1}\})\sqcup\{\mu\}=\{\alpha_{s},\mu,n-1\}$.
  1. (i)
    We have $\alpha_{s}=3$ and $s_{1}\in V(W_{n})\setminus(\{0,1,2,3,\ldots,2m-4\}\sqcup\{n-1\})$. Observe that $s_{1}\in\mathcal{U}$. Therefore, $F$ is a spanning facet by Proposition 3.12.
  2. (ii)
    We have $\alpha_{s}\in\{4,5,\ldots,m-2\}$ and $s_{1}\in V(W_{n})\setminus\{\alpha_{s}-4,\alpha_{s}-3,\alpha_{s}-2\}\sqcup\{\alpha_{s},\alpha_{s}+1,\ldots,2m-\alpha_{s}-1\}\sqcup\{n-1\}$.
    If $s_{1}\neq\alpha_{s}-1$, then $s_{1}\in\mathcal{U}$, and thus, $F$ is a spanning facet by Proposition 3.12.
    So assume that $s_{1}=\alpha_{s}-1$. We consider three cases: $\mu<\alpha_{s}$, $\mu\in\{\alpha_{s}+1,\alpha_{s}+2\}$, and $\mu>\alpha_{s}+2$.
    First, let $\mu<\alpha_{s}$. Then $\mu<s_{1}=\alpha_{s}-1$ (as $\mu\in F$). Moreover, $\alpha_{s}\leq m-2$ implies that $\alpha_{s}<_{\mathcal{O}}\mu$. Hence $F^{\prime\prime}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. Since $64\leq\alpha_{s}\leq m-2\leq n-$ 6, we have $\alpha_{s}\nsim n-1$. Further, $\mu<\alpha_{s}$ implies that either $\mu\in\{0,1\}$, or $2\leq\mu<\alpha_{s}$. Hence, $\mu\nsim\alpha_{s}$ or $\mu\nsim n-1$. Therefore $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\mu<s_{1}$, we have $F^{\prime\prime}\ll F$. Observe that $F,F^{\prime\prime}\notin\mathcal{D}$ (as $\alpha_{s}<m-1$). Therefore, $F^{\prime\prime}\prec F$ Definition 3.4 (i) and we take $\tilde{F}=F^{\prime\prime}$.
    Now, let $\mu\in\{\alpha_{s}+1,\alpha_{s}+2\}$. We have $14\leq\alpha_{s}<\mu\leq\alpha_{s}+2\leq m\leq n-4<n-$ 1. Thus, $\alpha_{s}\nsim n-1$ and $\mu\nsim n-1$. Hence $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{s}\leq m-2$, we have $\mu<_{\mathcal{O}}\alpha_{s}$. Therefore, $F^{\prime\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime\prime}$.
    Lastly, assume that $\mu>\alpha_{s}+2$. Then $13\leq\alpha_{s}-1=s_{1}<\alpha_{s}<\alpha_{s}+2<\mu<n-$ 1 implies that $\mu\nsim\alpha_{s}$ and $\mu\nsim s_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. Now, if $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\ll F$ by Proposition 3.11. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Hence, $F^{\prime}\prec F$ by Definition 3.4 (i). Therefore, we take $\tilde{F}=F^{\prime}$.
  3. (iii)
    We have $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{m-1,m,m+1,m+2,\ldots,n-3\}$, $s_{1}\in V(W_{n})\setminus(\{\omega,\omega+1,\ldots,\alpha_{s},\alpha_{s}+1,\alpha_{s}+2,\alpha_{s}+3\text{ (mod $n$})\}\sqcup\{y\})$, $\omega=\min\{2m-\alpha_{s},\alpha_{s}-4\}$ and
    $y=\begin{cases}n-1&\text{ if }\alpha_{s}<n-4,\\ 0&\text{ if }\alpha_{s}=n-4,\\ 1&\text{ if }\alpha_{s}>n-4.\end{cases}$
    First, suppose that $s_{1}\neq\alpha_{s}+4$ (mod $n$). Since $s_{1}\notin\{\omega,\omega+1,\ldots,\alpha_{s}+3\text{ (mod $n$)}\}$ and $\omega\leq\alpha_{s}-4$, it follows that $s_{1}<\alpha_{s}-4$ or $s_{1}>\alpha_{s}+4$ (mod $n$). Thus $|\alpha_{s}-s_{1}|\geq 5$. Now, since $\alpha_{s}\geq m-1>3$ and $s_{1}<n-1$, we get $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. We have $\alpha_{s}\nsim s_{1}$. Therefore $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. Now, if $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\ll F$ by Proposition 3.11. Since $s_{1}<\alpha_{s}-4$ or $s_{1}>\alpha_{s}+4$ (mod $n$), we have $F,F^{\prime}\notin\mathcal{D}$. Hence $F^{\prime}\prec F$ by Definition 3.4 (i). Therefore, we take $\tilde{F}=F^{\prime}$.
    Now, let $s_{1}=\alpha_{s}+4$ (mod $n$). Observe that if $\alpha_{s}\geq n-5$, then $\alpha_{s}+4\ \text{(mod $n$)}=y$. Therefore, since $s_{1}\neq y$, we have $\alpha_{s}\leq n-6$. This means that $\alpha_{s}+4\ \text{(mod $n$)}=\alpha_{s}+4$. First, let $\mu\notin\{\alpha_{s}+2,\alpha_{s}+3\}$. We have $14\leq m-1\leq\alpha_{s}<\alpha_{s}+4=s_{1}<n-$ 1. It follows that $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Further, since $\alpha_{s}\nsim s_{1}$, we get $F^{\prime}\in M(\Delta_{3}(W_{n}))$.
    If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. So assume that $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\ll F$ by Proposition 3.11. Since $n-1>s_{1}=\alpha_{s}+4$ and $\mu\neq\alpha_{s}+3$, we have $F,F^{\prime}\notin\mathcal{D}$. Therefore, $F^{\prime}\prec F$ by Definition 3.4 (i), and we take $\tilde{F}=F^{\prime}$.
    Now, let $\mu\in\{\alpha_{s}+2,\alpha_{s}+3\}$. Observe that since $\alpha_{s}\geq m-1$, we have $\alpha_{s}<_{\mathcal{O}}\mu$, and hence $F^{\prime\prime}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. We have $14\leq m-1\leq\alpha_{s}<\alpha_{s}+2<\alpha_{s}+3\leq n-3<n-$ 1 (as $\alpha_{s}\leq n-6$). Clearly, $\alpha_{s}\nsim n-1$. Moreover, if $\mu=\alpha_{s}+2$, then $\mu\nsim n-1$; and if $\mu=\alpha_{s}+3$, then $\mu\nsim\alpha_{s}$. Thus $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\mu<s_{1}$, we have $F^{\prime\prime}\ll F$. Further, $\alpha_{s}\leq n-6$ implies that $n-1\geq\alpha_{s}+5$ and thus, $F,F^{\prime\prime}\notin\mathcal{D}$. Hence, $F^{\prime\prime}\prec F$ by Definition 3.4 (i), and we take $\tilde{F}=F^{\prime\prime}$.
    Thus, in each case, we have a facet $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Thus, $F$ is a spanning facet.
Proposition 3.14.
Suppose $F\in M(\Delta_{3}(W_{n}))$ be such that either $F^{c}=\{x,x+1\ (\text{mod $n$}),x+4\ (\text{mod $n$})\}$, or $F^{c}=\{x,x+3\ (\text{mod $n$}),x+4\ (\text{mod $n$})\}$ for some $x\in V(W_{n})$. Then $F$ is not a spanning facet.
Proof.
We have $x+2$ (mod $n$) $\in F$. Let ${F^{\nu}}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{x+2\text{ (mod $n$)}\}$ for $\nu\in F^{c}$. Then for any $\nu\in F^{c}$, $W_{n}[{F^{\nu}}^{c}]$ is connected. Hence, there does not exist a facet $\tilde{F}$ such that $\tilde{F}\prec F$ and $\tilde{F}\cap F=F\setminus\{x+2\text{ (mod \ $n$)}\}$. Therefore, $F$ is not a spanning facet by definition. ∎
Remark 3.15.
Observe that if $F\in\mathcal{D}$, then either $F^{c}=\{x,x+1\ (\text{mod \ $n$}),x+4\ (\text{mod \ $n$})\}$ or $F^{c}=\{x,x+3\ (\text{mod \ $n$}),x+4\ (\text{mod \ $n$})\}$ for some $x\in V(W_{n})$. Therefore, if $F\in\mathcal{D}$, then $F$ is not a spanning facet by Proposition 3.14.
Proposition 3.16.
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$ for some $s\neq t$ such that ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and ${F^{\prime}}^{c}=\{\mu,s_{1},s_{2}\}$ for some $\mu\in F$. If $\alpha_{s}<_{\mathcal{O}}\mu$, then $F\prec F^{\prime}$.
Proof.
It follows that $F^{c}=(F^{\prime c}\setminus\{\mu\})\sqcup\{\alpha_{s}\}$ and since $F^{\prime}\in\mathcal{T}_{t}$, $\alpha_{t}\in\{\mu,s_{1},s_{2}\}$. Then, $\alpha_{s}<_{\mathcal{O}}\mu,s_{1},s_{2}$ implies that $F\prec F^{\prime}$ by Proposition 3.5. ∎
Proposition 3.17.
Let $F\in M(\Delta_{3}(W_{n}))$ be a spanning facet such that $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Then $\alpha_{s}\notin\{0,1,2,n-2,n-1\}$ and $s_{2}=n-1$.
Proof.
By Remark 3.3, we have $\alpha_{s}\notin\{0,1,n-1\}$. Now, since $F$ is a spanning facet, we get $\alpha_{s}\notin\{2,n-2\}$ by Proposition 3.14. Hence $\alpha_{s}\notin\{0,1,2,n-2,n-1\}$.
Observe that $\alpha_{s}<_{\mathcal{O}}n-1$. Suppose $n-1\notin F^{c}$. Then Remark 3.10 implies that there exists $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{n-1\}$ for some $\nu\in F^{c}$.
If $\nu=\alpha_{s}$, then by Proposition 3.16, $\alpha_{s}<_{\mathcal{O}}n-1$ implies that $F\prec\tilde{F}$, a contradiction. So, $\nu\in\{s_{1},s_{2}\}$. Since $\alpha_{s}<_{\mathcal{O}}s_{1},s_{2},n-1$ and $s_{1}<s_{2}<n-1$, we have either $\tilde{F}=\{\alpha_{s}\}\sqcup\{s_{2},n-1\}$ or $\tilde{F}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$. This implies that $F\in\mathcal{T}_{s}$ and $F\ll\tilde{F}$. Since $F$ is a spanning facet, $F\notin\mathcal{D}$ by Remark 3.15. Therefore, $F\prec\tilde{F}$ by Definition 3.4 (i) and (iii), again a contradiction. Thus, $n-1\in F^{c}$.
Clearly, $\alpha_{s}<_{\mathcal{O}}n-1$ and $s_{1}<s_{2}$ implies that $s_{2}=n-1$. ∎
Proposition 3.18.
Let $F\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ and ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Suppose there exists $\mu\in F$ such that $\alpha_{s}<_{\mathcal{O}}\mu$ and $s_{1}<\mu$. Let ${F^{\prime}}^{c}=({F}^{c}\setminus\{s_{2}\})\sqcup\{\mu\}$. If either $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, or $F^{\prime}\in\mathcal{D}$ whenever $F^{\prime}\in M(\Delta_{3}(W_{n}))$, then $F$ is not a spanning facet.
Proof.
Assume that either $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, or $F^{\prime}\in\mathcal{D}$ whenever $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Suppose $F$ is a spanning facet. We have $\mu\in F$. By Remark 3.10, there exists $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Moreover, $F\notin\mathcal{D}$ by Remark 3.15, and $s_{2}=n-1$ by Proposition 3.17.
If $\nu=\alpha_{s}$, then $\alpha_{s}<_{\mathcal{O}}\mu$ implies that $F\prec\tilde{F}$ by Proposition 3.16, a contradiction. Also, if $\nu=s_{1}$, then $\alpha_{s}<_{\mathcal{O}}\mu$ implies that $\tilde{F}^{c}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. It follows that $\tilde{F}\in\mathcal{T}_{s}$, and since $s_{1}<\mu$, $F\ll\tilde{F}$. Hence $F\notin\mathcal{D}$ implies that $F\prec\tilde{F}$ by Definition 3.4 (i) and (iii), again a contradiction. So, $\nu=s_{2}=n-1$. Then $\tilde{F}=F^{\prime}$.
If $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, then this implies that $\tilde{F}$ is not a facet, which is a contradiction.
If $F^{\prime}\in M(\Delta_{3}(W_{n}))$, then by our assumption, $F^{\prime}\in\mathcal{D}$. Now, since $\alpha_{s}<_{\mathcal{O}}\mu$ and $s_{1}<\mu$, we get $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s_{1},\mu\}$ and thus $F^{\prime}\in\mathcal{T}_{s}$. Further, since $F\notin\mathcal{D}$, $F\prec F^{\prime}$ by Definition 3.4 (iii), which is a contradiction.
Thus $F$ is not a spanning facet. ∎
Lemma 3.19.
Let $F\in M(\Delta_{3}(W_{n}))$ be a spanning facet. Then $F\in\mathcal{S}$.
Proof.
Let $F\in\mathcal{T}_{p}$ such that $F^{c}=\{\alpha_{p}\}\sqcup\{p_{1},p_{2}\}$. From Proposition 3.17, $\alpha_{p}\notin\{0,1,2,n-2,n-1\}$ and $p_{2}=n-1$. Hence $\alpha_{p}\in\{3,4,\ldots,n-3\}$. For each $\alpha_{p}$, we find the possible values of $p_{1}$ for which $F$ is a spanning facet and show that $F\in\mathcal{S}$ for all such values of $p_{1}$. We have $45\leq m\leq n-$ 4 (as $n\geq 9$). We consider the following five cases based on the values of $\alpha_{p}$.
  1. (i)
    $\alpha_{p}=3$.
    We show that $F$ satisfies $(\mathcal{S}_{1})$. For this, we need to show that $p_{1}\in V(W_{n})\setminus(\{0,1,2,3,4,\ldots,2m-4\}\sqcup\{n-1\})$. We have $\alpha_{p}=3<m$ (as $m\geq 5$) and $\alpha_{p}<_{\mathcal{O}}p_{1}$. By Remark 3.1 (i), either $p_{1}<3$ or $p_{1}\geq 2m-3$. Thus $p_{1}\notin\{3,4,\ldots,2m-4\}$.
    Suppose $p_{1}<3$. Then $W_{n}[F^{c}]$ is disconnected implies that $p_{1}\in\{0,2\}$. It follows that either $F^{c}=\{n-1,(n-1)+1\ (\text{mod \ $n$}),(n-1)+4\ (\text{mod \ $n$})\}$ or $F^{c}=\{n-1,(n-1)+3\ (\text{mod \ $n$}),(n-1)+4\ (\text{mod \ $n$})\}$, which is a contradiction to Proposition 3.14. Hence, $p_{1}\geq 2m-3$. Since $p_{1}<p_{2}=n-1$, $p_{1}\neq n-1$. This means that $p_{1}\in V(W_{n})\setminus(\{0,1,2,3,4,\ldots,2m-4\}\sqcup\{n-1\})$.
  2. (ii)
    $\alpha_{p}\in\{4,5,\ldots,m-2\}$.
    In this case, we show that $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2\}\sqcup\{\alpha_{p},\alpha_{p}+1,\ldots,2m-\alpha_{p}-1\}\sqcup\{n-1\})$, which implies that $F$ satisfies $(\mathcal{S}_{2})$.
    Since $\alpha_{p}\leq m-2$ and $\alpha_{p}<_{\mathcal{O}}p_{1}$, either $p_{1}<\alpha_{p}$ or $p_{1}\geq 2m-\alpha_{p}$ by Remark 3.1 (i). Thus $p_{1}\notin\{\alpha_{p},\alpha_{p}+1,\ldots,2m-\alpha_{p}-1\}$. Also, $p_{1}<p_{2}=n-1$ implies that $p_{1}\neq n-1$.
    Suppose $p_{1}\in\{\alpha_{p}-4,\alpha_{p}-3\}$. We have $p_{1}+2\leq\alpha_{p}-1\leq m-3<n-1=p_{2}$. Thus $p_{1}+2\in F$. Clearly, $p_{1}<p_{1}+2$. Moreover, observe that $\alpha_{p}<_{\mathcal{O}}p_{1}+2$, and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+2\}]=W_{n}[\{p_{1},p_{1}+2,\alpha_{p}\}]$ is connected. Thus $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, where $F^{\prime c}=(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+2\}$. It follows from Proposition 3.18 that $F$ is not a spanning facet, a contradiction. Therefore $p_{1}\notin\{\alpha_{p}-4,\alpha_{p}-3\}$.
    Now, suppose $p_{1}=\alpha_{p}-2$. Since $W_{n}[F^{c}]$ is disconnected, $p+1=\alpha_{p}-1\in F$. We have $\alpha_{p}<_{\mathcal{O}}\alpha_{p}-1=p+1$ and $p_{1}<p_{1}+1$. Moreover, $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}]=W_{n}[\{\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}\}]$ is connected. Hence, using Proposition 3.18, $F$ is not a spanning facet, a contradiction. Thus, $p_{1}\neq\alpha_{p}-2$.
    Therefore, $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p},\alpha_{p}+1,\ldots,2m-\alpha_{p}-1\}\sqcup\{n-1\})$.
  3. (iii)
    $\alpha_{p}\in\{m-1,m\}$.
    Let
    $y=\begin{cases}n-1&\text{ if }\alpha_{p}<n-4,\\ 0&\text{ if }\alpha_{p}=n-4,\\ 1&\text{ if }\alpha_{p}>n-4.\end{cases}$
    We show that $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$, which implies that $F$ satisfies $(\mathcal{S}_{3})$ (as $\min\{2m-\alpha_{p},\alpha_{p}-4\}=\alpha_{p}-4$).
    We have $F^{c}=\{\alpha_{p}\}\sqcup\{p_{1},p_{2}\}$, where $p_{2}=n-1$. Clearly, $p_{1}\notin\{\alpha_{p},n-1\}$.
    Suppose $p_{1}=y$. We have $\alpha_{p}\leq m\leq n-4$ (as $n\geq 9$). So $y\neq 1$. Since $p_{1}\neq n-1,$ we get $y=0$. Clearly, if $\alpha_{p}\neq n-4$, then $y\neq 0$. Thus $\alpha_{p}=n-4$. This means that $n=9$ and $\alpha_{p}=m$, thereby implying that $F^{c}=\{m,m+3,m+4\ (\text{mod $n$)}\}$. Then $F$ is not a spanning facet by Proposition 3.14. This is a contradiction. So $p_{1}\neq y$.
    We now show that if $p_{1}$$\in\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}+1,\alpha_{p}+2\}$, then there exists $\mu\in F$ such that $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]$ is connected. Then Proposition 3.18 implies that $F$ is not a spanning facet, which is a contradiction.
    1. (a)
      $p_{1}\in\{\alpha_{p}-4,\alpha_{p}-3\}$.
      We have $p_{1}+2\leq\alpha_{p}-1\leq m-1<n-1=p_{2}$. It follows that $p_{1}+2\in F$. Take, $\mu=p_{1}+2$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{p_{1},p_{1}+2,\alpha_{p}\}]$ is connected.
    2. (b)
      $p_{1}=\alpha_{p}-2$.
      Since $p_{1}+1=\alpha_{p}-1\leq m-1<n-1=p_{2}$, we have $p_{1}+1\in F$. Take $\mu=p_{1}+1$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}\}]$ is connected.
    3. (c)
      $p_{1}=\alpha_{p}-1$.
      Here, $p_{1}+3=\alpha_{p}+2\leq m+2<n-1=p_{2}$, which implies that $p_{1}+3\in F$. Take $\mu=p_{1}+3$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p}-1,\alpha_{p},\alpha_{p}+2\}]$ is connected.
    4. (d)
      $p_{1}=\alpha_{p}+1$.
      In this case, $\alpha_{p}<_{\mathcal{O}}p_{1}$ implies that $\alpha_{p}=m$. We have $p_{1}+1=m+2<n-1$. So $p_{1}+1\in F$. Take, $\mu=p_{1}+1$. We have $\alpha_{p}=m<_{\mathcal{O}}m+2=p_{1}+1=\mu$ and $p_{1}<\mu$. Clearly, $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p},\alpha_{p}+1,\alpha_{p}+2\}]$ is connected.
    5. (e)
      $p_{1}=\alpha_{p}+2$.
      Here, since $W_{n}[F^{c}]$ is disconnected, we have $n\geq 10$. It follows that $p_{1}+1=\alpha_{p}+3\leq m+3<n-1$, and thus $p_{1}+1\in F$. Take $\mu=p_{1}+1$. Observe that $\alpha_{p}<_{\mathcal{O}}\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p},\alpha_{p}+2,\alpha_{p}+3\}]$ is connected.
    Thus, $p_{1}\notin\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2\}\sqcup\{y\}$.
    Now, suppose $p_{1}=\alpha_{p}+3\ (\text{mod \ $n$})$. Since $\alpha_{p}\leq m$, we have $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. We first assume that $n=9$. Then, $m+3=n-1$. Therefore, if $\alpha_{p}=m-1$, then $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\}$, which implies that $F$ is not a spanning facet by Proposition 3.14. This is a contradiction. So $\alpha_{p}=m$. Then $p_{1}=m+3=n-1=p_{2}$, which is again a contradiction.
    Thus, we assume that $n\geq 10$. It follows that $p_{1}+1=\alpha_{p}+4\leq m+4\leq n-1$. If $p_{1}+1=n-1$, then $\alpha_{p}=m$ and $F^{c}=\{m,m+3,m+4\}$. By Proposition 3.14, it follows that $F$ is not a spanning facet, a contradiction. Hence $p_{1}+1<n-1$ and thus, $p_{1}+1\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. Let $F^{\prime c}=(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}$. If $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, then Proposition 3.18 implies that $F$ is not a spanning facet, which is a contradiction. So $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Then $F^{\prime c}=\{\alpha_{p}\}\sqcup\{\alpha_{p}+3,\alpha_{p}+4\}$ satisfies $(\mathcal{D}_{4})$, and thus $F^{\prime}\in\mathcal{D}$. Hence, Proposition 3.18 implies that $F$ is not a spanning facet, again a contradiction. So $p_{1}\neq\alpha_{p}+3\text{ (mod \ $n$})$.
    Therefore, $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
  4. (iv)
    $\alpha_{p}=m+1$.
    Let
    $y:=\begin{cases}n-1&\text{ if }m+1<n-4,\\ 0&\text{ if }m+1=n-4,\\ 1&\text{ if }m+1>n-4.\end{cases}$
    We show that $F$ satisfies $(\mathcal{S}_{3})$. For this, we show that $p_{1}\in V(W_{n})\setminus(\{m-3,m-2,m-1,m,m+1,m+2,m+3,m+4\ (\text{mod $n$})\}\sqcup\{y\})$ (as $\min\{2m-\alpha_{p},\alpha_{p}-4\}=\alpha_{p}-4=m-3$). We have $F^{c}=\{m+1\}\sqcup\{p_{1},p_{2}\}$, where $p_{2}=n-1$.
    Suppose $p_{1}=y$. Since $p_{1}\neq p_{2}=n-1$, we get $y\in\{0,1\}$. If $y=0$, then $m+1=n-4$. Hence $p_{2}=n-1=m+4$, which implies that $F^{c}=\{m+1,(m+1)+3\text{ (mod \ $n$}),(m+1)+4\text{ (mod \ $n$})\}$. Thus, $F$ is not a spanning facet by Proposition 3.14. This is a contradiction. So $y=1$, and then $m+1>n-4$. This implies that $W_{n}[F^{c}]$ is connected, again a contradiction. Hence $p_{1}\neq y$.
    Since $\alpha_{p}=m+1<_{\mathcal{O}}p_{1}$, we have $p_{1}\notin\{m-1,m,m+1\}$. Suppose $p_{1}\in\{m-3,m-2,m+2,m+3,m+4\ (\text{mod $n$})\}$.
    1. (a)
      $p_{1}=m-3$.
      Here, $p_{1}+1=m-2\in F$. We have $\alpha_{p}=m+1<_{\mathcal{O}}m-2=p_{1}+1$ and $p_{1}<p_{1}+1$. Observe that if $F^{\prime c}=\{m+1\}\sqcup\{m-3,m-2\}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ such that $F^{\prime}$ satisfies $(\mathcal{D}_{3})$ and thus, $F^{\prime}\in\mathcal{D}$. Therefore, Proposition 3.18 implies that $F$ is not a spanning facet, a contradiction. Hence $p_{1}\neq m-3$.
    2. (b)
      $p_{1}=m-2$.
      We have $p_{1}+4=m+2\in F$. Also, $\alpha_{p}=m+1<_{\mathcal{O}}m+2=p_{1}+4$ and $p_{1}<p_{1}+4$. If $F^{\prime c}=\{m+1\}\sqcup\{m-2,m+2\}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ such that $F^{\prime}$ satisfies $(\mathcal{D}_{1})$. Thus $F^{\prime}\in\mathcal{D}$. Therefore, by Proposition 3.18, $F$ is not a spanning facet. This is a contradiction. Hence $p_{1}\neq m-2$.
    3. (c)
      $p_{1}\in\{m+2,m+3\}$.
      Observe that if $n=9$, then $n-1=m+3$; if $n\in\{10,11\}$, then $n-1=m+4$; and if $n\in\{12,13\}$, then $n-1=m+5$. We have $F^{c}=\{m+1\}\sqcup\{p_{1},n-1\}$. Since $W_{n}[F^{c}]$ is disconnected and $p_{1}<n-1$, it follows that if $p_{1}=m+2$, then $n\geq 12$; and if $p_{1}=m+3$, then $n\geq 14$. Therefore, in each case, we have $p_{1}+2<n-1$. This implies that $p_{1}+2\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+2$, $p_{1}<p_{1}+2$ and $W_{n}[(F^{c}\setminus\{n-1\})\sqcup\{p_{1}+2\}]=W_{n}[\{\alpha_{p},p_{1},p_{1}+2\}]$ is connected. Using Proposition 3.18, we get a contradiction to the fact that $F$ is a spanning facet. Hence $p_{1}\notin\{m+2,m+3\}$.
    4. (d)
      $p_{1}=m+4\ (\text{mod $n$})$.
      Then $F^{c}=\{m+1,m+4\ (\text{mod $n$}),n-1\}$. Since $W_{n}[F^{c}]$ is disconnected and $p_{1}<n-1$, we get $n\geq 12$. Further, if $n\in\{12,13\}$, then $n-1=m+5$, and thus, $F^{c}=\{m+1,(m+1)+3,(m+1)+4\}$. It follows from Proposition 3.14 that $F$ is not a spanning facet. This is a contradiction. So, assume that $n\geq 14$. Then $p_{1}=m+4\ (\text{mod $n$})=m+4$ and $p_{1}+1=m+5<n-1$. Thus $p_{1}+1\in F$. Let $F^{\prime c}=\{m+1\}\sqcup\{m+4,m+5\}$. Then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ and satisfies $(\mathcal{D}_{4})$. Hence $F^{\prime}\in\mathcal{D}$. Using Proposition 3.18, it follows that $F$ is not a spanning facet, a contradiction. So $p_{1}\neq m+4\ (\text{mod $n$})$
    Therefore $p_{1}\not\in\{m-3,m-2,m-1,m,m+1,m+2,m+3,m+4\ (\text{mod $n$})\}\sqcup\{y\}$.
  5. (v)
    $\alpha_{p}\in\{m+2,\ldots,n-3\}$.
    We show that $F$ satisfies $(\mathcal{S}_{3})$. Since $\min\{2m-\alpha_{p},\alpha_{p}-4\}=2m-\alpha_{p}$, we prove that $p_{1}\in V(W_{n})\setminus(\{2m-\alpha_{p},\ldots,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\text{ (mod \ $n$})\}\sqcup\{y\})$, where
    $y=\begin{cases}n-1&\text{ if }\alpha_{p}<n-4,\\ 0&\text{ if }\alpha_{p}=n-4,\\ 1&\text{ if }\alpha_{p}>n-4.\end{cases}$
    We have $\alpha_{p}\geq m+2$ and $\alpha_{p}<_{\mathcal{O}}p_{1}$. Thus, either $p_{1}<2m-\alpha_{p}$ or $p_{1}>\alpha_{p}$ by Remark 3.1 (ii). This means that $p_{1}\notin\{2m-\alpha_{p},\ldots,\alpha_{p}-1,\alpha_{p}\}$. Therefore, it is sufficient to show that $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\text{ (mod \ $n$})\}\sqcup\{y\}$.
    1. (a)
      $\alpha_{p}\in\{m+2,m+3,\ldots,n-6\}$.
      Here, since $\alpha_{p}\leq n-6<n-4$, we have $y=n-1$. Clearly, $p_{1}\neq y$. Suppose $p_{1}\in\{\alpha_{p}+1,\alpha_{p}+2\}$. Then $p_{1}+1\leq\alpha_{p}+3\leq n-3<n-1$, and thus, $p_{1}+1\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. Further, since $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}]=W_{n}[\{\alpha_{p},p_{1},p_{1}+1\}]$ is connected, $F$ is not a spanning facet by Proposition 3.18. This is a contradiction. Therefore $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2\}$.
      Suppose $p_{1}=\alpha_{p}+3\text{ (mod \ $n$})$. We have $\alpha_{p}+3\leq n-3<n-1$. So $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. Then $p_{1}+1=\alpha_{p}+4\leq n-2<n-1$, and thus, $p_{1}+1\in F$. Clearly, $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. If $F^{\prime c}=\{\alpha_{p}\}\sqcup\{\alpha_{p}+3,\alpha_{p}+4\}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ and satisfies $(\mathcal{D}_{4})$. Thus $F^{\prime}\in\mathcal{D}$. By Proposition 3.18, it follows that $F$ is not a spanning facet, a contradiction. Therefore, $p_{1}\neq\alpha_{p}+3\text{ (mod \ $n$})$.
    2. (b)
      $\alpha_{p}=n-5$.
      We have $y=n-1$ and $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. Further, $\alpha_{p}+4=n-1=p_{2}$. Therefore, since $W_{n}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, it follows that $p_{1}\notin\{\alpha_{p}+2,y\}$. Moreover, if $p_{1}\in\{\alpha_{p}+1,\alpha_{p}+3\}$, then either $F^{c}=\{\alpha_{p},\alpha_{p}+1,\alpha_{p}+4\}$ or $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\}$. Then, by Proposition 3.14, it follows that $F$ is not a spanning facet, a contradiction. Hence $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+3\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
    3. (c)
      $\alpha_{p}=n-4$.
      Then $y=0$ and $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3=n-1=p_{2}$. If $p_{1}=y$, then $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\ (\text{mod \ $n$})\}$, which contradicts the fact that $F$ is a spanning facet by Proposition 3.14. Hence $p_{1}\neq y$. Further, since $W_{n}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, we get $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
    4. (d)
      $\alpha_{p}=n-3$.
      We have $y=1$. Since $W_{n}[F^{c}]$ is disconnected and $p_{1}<p_{2}$, it follows that $p_{1}\notin\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$}),1)\}$. Therefore, $p_{1}\not\in\{\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
This completes the proof of Lemma 3.19. ∎
Proof of Theorem 1.4.
In Section 3.1, we have proved that $\prec$ gives a shelling order for $\Delta_{3}(W_{n})$. Hence the $3$-cut complex $\Delta_{3}(W_{n})$ is shellable for $n\geq 9$.
Now, using Lemmas 3.13 and 3.19, we see that $F\in M(\Delta_{3}(W_{n}))$ is a spanning facet for the shelling order given by $\prec$ on $\Delta_{3}(W_{n})$ if and only if $F\in\mathcal{S}$. Therefore, to count the number of spanning facets, it is enough to count the total facets in $\mathcal{S}$. Let $F\in\mathcal{S}$. Then $F\in\mathcal{T}_{s}$ for some $s$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$ and $F$ satisfies one of the conditions from $(\mathcal{S}_{1})$ to $(\mathcal{S}_{3})$.
If $F$ satisfies $(\mathcal{S}_{1})$, then $\alpha_{s}=3$ and there are $n-(2m-2)=n-2m+2$ possibilities for $s_{1}$ and hence the same number of spanning facets satisfying $(\mathcal{S}_{1})$.
If $F$ satisfies $(\mathcal{S}_{2})$, then there are $m-5$ possibilities for $\alpha_{s}$, and for each value of $\alpha_{s}$, there are $n-(2m-2\alpha_{s}+4)$ possibilities for $s_{1}$. Therefore, the number of spanning facets satisfying $(\mathcal{S}_{2})$
$=(m-5)(n-2m-4)+\left(2\sum_{\alpha_{s}=4}^{m-2}\alpha_{s}\right)$
$=(m-5)(n-2m-4)+(m-2)(m-1)-12$
$=mn-m^{2}+3m-5n+10.$
Suppose $F$ satisfies $(\mathcal{S}_{3})$. If $\alpha_{s}\in\{m-1,m,m+1\}$, then $\omega=\alpha_{s}-4$. For each $\alpha_{s}$, there are $n-9$ possibilities for $s_{1}$. Therefore, we have $3(n-9)$ spanning facets. If $\alpha_{s}\in\{m+2,\ldots,n-3\}$, then $\omega=2m-\alpha_{s}$. There are $n-m-4$ possibilities for $\alpha_{s}$, and for each value of $\alpha_{s}$, there are $n-(2\alpha_{s}-2m+5)$ possibilities for $s_{1}$. Thus, the number of spanning facets satisfying $(\mathcal{S}_{3})$
$=(n-m-4)(n+2m-5)-\left(2\sum_{\alpha_{s}=m+2}^{n-3}\alpha_{s}\right)$
$=(n-m-4)(n+2m-5)-(n-3)(n-2)+(m+1)(m+2)$
$=mn-m^{2}-4n+16.$
Therefore, the total number of spanning facets
$=(n-2m+2)+(mn-m^{2}+3m-5n+10)+3(n-9)$
$\quad+(mn-m^{2}-4n+16)$
$=2mn-2m^{2}+m-5n+1$
$=\frac{n^{2}-9n+2}{2}$
$=\binom{n-4}{2}-9.$
Hence from Theorem 2.6, we get $\Delta_{3}(W_{n})\simeq\bigvee\limits_{\binom{n-4}{2}-9}\mathbb{S}^{n-4}.$
Acknowledgement
The first author is supported by HTRA fellowship by IIT Mandi, India. The second author is supported by the seed grant project IITM/SG/SMS/95 by IIT Mandi, India. The third author is partially supported by the DISHA fellowship by NISER Bhubaneswar, India.